### v1

# Design principles of the cytotoxic CD8+ T-cell response

Abstract
Cytotoxic T lymphocytes eliminate infected or malignant cells, safeguarding surrounding tissues. Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive. Here, we recast a T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues. We demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings. We recover an inherent trade-off: infection clearance at the cost of immunopathology. We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals. This model explains experimentally observed patterns of T-cell effector expansion regulation in mice. Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment. Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells. Our findings offer a unified control-logic for cytotoxic T-cell responses and point to specific regulatory programs that can be engineered for more robust T-cell therapies.
I Introduction
Cellular immunity hinges upon the regulated activation, expansion, and contraction of lymphocytes that clear an infection while limiting collateral tissue damage. We focus on CD8+ T-cells—also known as killer or cytotoxic T-cells—which are crucial for immune surveillance and defense against infections and cancer. Pathogen recognition is mediated by T-cell receptors (TCRs) that bind short pathogen-derived peptides (or antigens) presented by major histocompatibility complexes (MHCs) on antigen-presenting cells (APCs) and other target cells. To recognize a multitude of pathogens, the immune system generates a diverse repertoire of TCRs through a stochastic process known as V(D)J recombination, producing hundreds of billions of distinct T-cell clonotypes (i.e., T-cells with unique receptors) [61, 46, 56].
Pathogen recognition triggers a response program in naïve T-cells, tuned by antigen signal, pro- and anti-inflammatory signaling molecules (cytokines), and other environmental signals generated by the infection and the host’s response (Fig. 1A) [72, 37, 49, 48, 21]. These signals determine the extent of burst-like proliferation of short-lived effector T-cells versus differentiation into long-lived, slowly dividing memory subsets. Effector T-cells have the capacity to rapidly divide—up to 19 cell divisions [8, 80, 13, 48]—and secrete cytolytic chemical molecules that can degrade infected or unwanted cells [63, 35]. While most activated naïve cells differentiate into terminal effectors that undergo apoptosis once infection is cleared, a small fraction survive as long-lived effector or central memory cells [66, 63, 78, 1].
The logic of signal integration for immune decision-making throughout a response remains the central open problem. Simple additive models of antigen and cytokine inputs can reproduce experimental patterns of clonal expansion in some settings [49, 48, 21]. However, genetic and single-cell analyses reveal far richer control circuits governing activation, differentiation, and contraction of T-cells that involve intricate genetic programs and probabilistic regulatory rules [32, 16, 1, 60, 77, 78, 13, 40].
Pathogens vary in their immunogenicity—the ease with which they are recognized as “non-self”—and in the degree of harm they impose via secreted toxins, inflammation, and disruption of cellular function. The CD8+ T-cell response must have the potential to curb “harm” across diverse infection scenarios [20, 52], yet the evolutionary objectives and biophysical constraints shaping their decision-making program remain unclear. Here, we formalize one plausible objective for T-cell decision-making: to minimize cumulative host healthy-cell loss due to infection and collateral death of healthy cells due to the immune response (immunopathology or toxicity).
We model the T-cell decision-making program as a flexible, coarse-grained signal-integration protocol. Specifically, we cast each CD8+ T-cell as a feedback controller whose intra-cellular regulatory program modulates activation, differentiation, proliferation, and death in response to intra-host cues, such as antigen levels and cytokine signals, which themselves are dynamic in the course of an infection [70, 48].
We then ask how alternative immunological objectives give rise to distinct effector programs. First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules. Next, by systematically sampling designs, we optimize for rapid pathogen clearance while minimizing immunopathology. We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology. Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [28, 41, 17, 54], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses. Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms. Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments—and thus candidate genetic engineering—that could enhance T-cell-based cancer immunotherapy.
The immune system’s true performance objective, and how evolution has fine-tuned it, may ultimately be unknowable. Formulating and testing principled heuristics provides a baseline to gauge the performance of real immune systems in responding to infections. This perspective has yielded deep insights in other biological contexts, revealing the architectural logic of genetic and regulatory networks [23, 67, 42, 31, 6]. In immunology, high-level mathematical models have demonstrated how a finite lymphocyte repertoire can be distributed to best confront a diverse pathogenic landscape [59, 50, 11] and how immune memory should be allocated to counter future threats [51, 64, 15, 65]. By extending this perspective to the dynamics of CD8+ T-cell signal integration and response, our work provides a quantitative framework that links observed decision-making to operating principles and trade-offs.
Figure 1: CD8+ T-cell response to infections. (A) The schematic shows how upon infection, susceptible cells $S$ become infected II at rate $b_{I}$. Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$. In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates. Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory. Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$). Infection dynamics $(S,I)$ are modeled by ODEs coupled to a stochastic dynamics of the T-cell response. Memory transitions (gray arrows) are neglected in our model. (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time. (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs. 6, 7). The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq. 8). (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq. 2) and immunogenicities $\kappa$ (eq. 3). For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm). Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq. 9). Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
Model
To characterize the immune response to an infection, we model the large-population dynamics of susceptible and infected cells in a host as deterministic processes, coupled to a stochastic, time-inhomogeneous birth-death process for the CD8+ T-cell response (Fig. 1A, and SI Appendix for details).
Intra-host infection dynamics.
We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Fig. 1A). The dynamics of the susceptible and infected cells follow,
$\begin{split}\dot{S}&=-b_{I}IS-d_{E}\frac{E}{K_{S}+I\frac{K_{S}+E}{K_{I}+E}+E+S}S\\ \dot{I}&=b_{I}IS-d_{I}I-d_{E}\frac{E}{K_{I}+S\frac{K_{I}+E}{K_{S}+E}+E+I}I,\end{split}$ (1)
where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are inverse affinity constants quantifying how well effector cells bind to and kill infected and susceptible cells, respectively [14, 25, 30]. The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix, section 1).
We quantify within-host reproductive success of a pathogen by the intra-host basic reproduction number,
$R_{0}=\frac{b_{I}S_{\text{max}}}{d_{I}},$ (2)
measuring the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population. This definition mirrors the standard epidemiological reproduction number [55, 2]. We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells. These inverse affinity parameters determine the immunogenicity of an infection, which we define as,
$\text{immunogenicity:}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \kappa=K_{I}^{-1}K_{S}.$ (3)
This dimensionless quantity reflects how readily cognate T-cells recognize antigens presented by infected cells as “non-self” relative to those presented by healthy cells (self). We consider infection scenarios with broad ranges of immunogenicity ($\kappa$) and basic reproduction numbers ($R_{0}$) (SI Appendix, section 1).
Signal cues to modulate T-cell decision-making.
We express damage during an infection in terms of cell deaths,
$\begin{split}h(t)&=h_{\text{inf}}(t)+\underbrace{h_{E,I}(t)+h_{E,S}(t)}_{=h_{\text{res}}(t)},\end{split}$ (4)
where $h_{\text{inf}}(t)$ represents the number of cells killed directly by the pathogen (infection), and $h_{E,I}(t)$ and $h_{E,S}(t)$ denote the number of infected and healthy cells, respectively, killed by the effector T-cell response, amounting to the total damage by the response $h_{\text{res}}(t)$.
[…]
We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$. Together, expressed by their inverse affinity constants, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
$\begin{split}&\text{antigen signal:}\penalty 10000\ \penalty 10000\ \sigma_{\text{Ag}}(t)={K_{I}}^{-1}{I}(t)\\ &\text{harm signals:}\penalty 10000\ \penalty 10000\ \sigma_{\text{inf}/\text{res}}(t)=K_{H}^{-1}\int_{0}^{t}e^{-(t-s)d_{H}}\,h_{\text{inf}/\text{res}}(s)\,{\rm d}s.\end{split}$ (5)
The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix, section 1).
By disentangling these sources of (potentially redundant) signals, we probe how CD8+ T-cells use these cues to make fate decisions. We do not pre-assign any of the signals to a pro- or anti-inflammatory role but allow optimization goals to determine their roles across diverse infection scenarios.
Design of T-cell programs for infection response.
Naïve T-cells cycle through lymphoid tissues, becoming activated when they durably engage cognate peptide-MHC complexes on activated APCs [71, 24]. Activated T-cells can differentiate into an effector phenotype, which participates in infection clearance, or into a long-lived memory state (Fig. 1A). Here, we examine the course of CD8+ T-cells response during an infection, dividing the process into four canonical response modules (Figure. 1A):
  1. 1.
    Activation: priming of naïve T-cells, $N\to N^{*}$,
  2. 2.
    Differentiation: commitment of the activated T-cell to an effector phenotype, $N^{*}\to E$,
  3. 3.
    Expansion: division of effector cells, $E\to 2E$,
  4. 4.
    Contraction: programmed loss of effector cells, $E\to\emptyset$.
These response processes map onto partially distinct genetic programs: the MYC pathway drives activation and proliferation [32]; the Wnt-Tcf7 pathway regulates effector versus memory fate decisions [26, 1]; and the Bcl-2-family of proteins control contraction of the effector cells via apoptosis [32].
For each process $i$, we represent its genetic regulation by a time-dependent transition rate $r_{i}(t)$, modulated by the instantaneous signals $\vec{\sigma}(t)$ (eq. 5) received by the T-cell (Fig. 1C),
$r_{i}(t)=$ $r^{\text{max}}\ g_{i}\left(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell_{i}\right),$ (6)
where the maximal rate $r^{\text{max}}$ is capped by the cell-cycle time [35, 40, 71], and $g_{i}$ is a monotonic response function of signal input, chosen from the modified Monod-Wyman-Changeux functions (Hill-type family) [19, 75]:
$\begin{split}&g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi}_{i},\ell_{i})=\frac{1}{1+\exp\left[-L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})\right]},\\ &\text{with,}\\ &L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})=\ell_{i}+\sum_{\alpha\in\{\text{Ag},\text{inf},\text{res}\}}\psi^{\alpha}_{i}\log\left[1+{\sigma_{\alpha}}\right].\end{split}$ (7)
The parameter $\ell_{i}$ sets the baseline propensity of transition $i$: in the absence of external signals, the regulatory factor evaluates to $g_{0,i}=(1+\exp(-\ell_{i}))^{-1}$ (Figure. 1C). The weights $\{\psi_{i}^{Ag},\psi_{i}^{\text{inf}},\psi_{i}^{\text{res}}\}$ encode the sensitivity and polarity (positive / negative = up /down-regulation) of response to the antigen ($\sigma_{\text{Ag}}$) and harm signals ($\sigma_{\text{inf}}$ and $\sigma_{\text{res}}$). This formulation abstracts away biochemical details while implementing the fold-change response to stimuli observed in many biological circuits [31, 58, 3, 74, 4].
To reflect shared regulatory programs while controlling complexity, we impose two simplifications. First, activation and early expansion are both MYC-driven [33]; we therefore require the modulatory factors, $g_{N\to N^{*}}$ and $g_{E\to 2E}$, to share parameter values, even though their realized rates will differ because signals vary in time, as cells transition between naive and effector states. Second, logically, since extracellular cues should converge on coherent effector programs, we enforce a common set of signal sensitivities across modules. With the sign convention that positive sensitivities promote effector production, for each signal $\sigma$, we set $\psi_{\sigma}\equiv\psi^{\sigma}_{N\to N^{*}}=\psi^{\sigma}_{E\to 2E}=\psi^{\sigma}_{N^{*}\to E}=-\psi^{\sigma}_{E\to\emptyset}$.
Figure 2: Design parameters shape T-cell response and infection clearance. (A) Varying the design parameters influences effector trajectories, summarized by four features: activation time, expansion rate, peak effector magnitude, and contraction rate. (B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom)—measured in units of maximum number of susceptible cells $S_{\text{max}}$—for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel. All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$. Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig. S1. (C) Sensitivities of trajectory (as in B) fold expansion, defined as the maximum effector count divided by the initial number of T-cells $N_{\text{lin.}}$, to changes in baseline transition rates (colors). Parameters are the same as Fig. 1, except for $K_{I}=10^{-2}S_{\text{max}}$.
The parameters of these constrained response functions define a response design $\vec{\theta}$ as,
$\vec{\theta}=(\underbrace{\psi_{\text{Ag}},\psi_{\text{inf}},\psi_{\text{res}}}_{\text{signal feedback}},\underbrace{\ell_{N\to N^{*}}=\ell_{E\to 2E},\ell_{N^{*}\to E},\ell_{E\to\emptyset}}_{\text{module-specific baselines}}).$ (8)
We systematically sample $\vec{\theta}$’s to assess performance of designs against different infections (Fig. 1D; SI Appendix, section 3). While $\vec{\theta}$ is broadly sampled, we hold the remaining biological rates and network parameters underlying the decision-making program in Fig. 1A within empirically supported ranges reported in SI Appendix, Table S1.
Cell economics of T-cell response.
Tracking all clinical sequelae of an infection is rarely feasible, but tissue integrity degrades as cells are lost. Cumulative cell death correlates with macroscopic readouts such as weight change, pathogen burden, and temperature dynamics [29, 43]. We therefore adopt a cell-economics objective in which harm is measured by the total number of cells lost over the course of infection.
We evaluate the performance of a given design $\vec{\theta}$ by simulating the response dynamics (Fig. 1D) and computing the total cell death (or inflicted harm) as
$\begin{split}H(\vec{\theta})=\underbrace{\int_{0}^{T}\left[h_{\text{inf}}\left(t,\vec{\theta}\right)+h_{E,I}\left(t,\vec{\theta}\right)\right]{\rm d}t}_{H_{\rm inf.}(\vec{\theta})}+\underbrace{\int_{0}^{T}h_{E,S}\left(t,\vec{\theta}\right){\rm d}t}_{H_{\rm toxic.}(\vec{\theta})},\end{split}$ (9)
where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, section 1 and 2, and Table S1). This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig. 1D).
Figure 3: Toxicity-infection harm trade-off across T-cell regulatory designs. (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq. 9)—measured in units of $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)). Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq. 8). Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed. Black points denote no-feedback designs ($\vec{\psi}=0$), and the large magenta circle in each panel marks the design that minimizes total harm, $\vec{\theta}^{(\mathrm{min})}$. (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A). By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I, II, III, IV (SI Appendix, section 3). (C) Toxicity vs. infection harm associated with optimal designs for each archetype is shown. Large dots represent cluster averages and small transparent dots represent individual infections, sampled in (B). (D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for archetype $i$ to within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)). Error bars denote the standard deviation within this ensemble. All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$). Unless otherwise noted, simulation parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$
Results
Cell-fate transitions determine effector response and pathogen clearance
How do rates of cell-fate transitions, determined by the regulatory design $\vec{\theta}$ (eq. 8), impact the immune response and pathogen clearance? To isolate the effect of baseline propensities ($\vec{\ell}$) independently of signal feedback ($\vec{\psi}$), we first simulated “hard-wired” designs in which all signal sensitivities were set to zero ($\vec{\psi}_{i}=0$). Each transition then happens at a fixed fraction of its maximal rate $g_{0,i}=[1+\exp(-\ell_{i})]^{-1}$, independently of antigen or harm signals; here, $i\in\{N\to N^{*}/E\to 2E,N^{*}\to E,E\to\emptyset\}$ indicates the transition module.
Varying these baseline rates produced strikingly different effector trajectories (Fig. 2). Promoting activation/proliferation ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\uparrow$), or lowering the contraction baseline rate ($g_{0,E\to\emptyset}\,\downarrow$) generated a rapid, high-amplitude effector burst that cleared the pathogen quickly, after which effector numbers decay (Fig. 2B, C). In contrast, slowing activation and division ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\downarrow$), or accelerating contraction ($g_{0,E\to\emptyset}\,\uparrow$), yielded sluggish effector accumulation, delayed clearance, and in extreme cases pathogen persistence, accompanied by a prolonged plateau of few effectors. Interestingly, varying the baseline rate of differentiation $g_{0,N^{*}\to E}$ had a weaker impact on the overall effector profile and pathogen clearance (Fig. 2B, C), suggesting relevance more downstream in the response program.
Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated cytokine signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig. S1).
Effector designs trade infection harm for toxicity
Previous research has identified a trade-off between maximizing infection clearance and minimizing response toxicity (immunopathology) in the immune response [36, 17, 54]. Consistently, by systematically sampling the space of designs (SI Appendix, section 3), we found this trade-off in the form of a Pareto front: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig. 3A). The extent of this trade-off depends on pathogen characteristics. This trade-off is stronger for infections with a higher intra-host reproductive number $R_{0}$ or lower immunogenicity $\kappa$. Interestingly, for scenarios with lower $R_{0}$, many of the hard-wired networks without signal feedback ($\vec{\psi}_{i}=0$) lie along the Pareto front, and perform near-optimally. For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far inside the sub-optimal region. Proportionate control in these cases requires designs with appropriate sensitivities to antigen and cytokine/harm signals ($\vec{\psi}\neq 0$) to minimize infection harm while limiting immunopathology.
Figure 4: Pareto-optimal designs balancing harm from acute infections and autoimmunity. (A) Bi-objective optimization to minimize the weighted sum of harms associated with T-cell regulatory designs $\vec{\theta}$ across two infection archetypes ($A,B$) yields a Pareto front (magenta curve); eq. 10. Points on the curve are Pareto-optimal designs, parameterized by the weight $\lambda\in[0,1]$, quantifying the emphasis on archetype B; moving left to right corresponds to decreasing $\lambda$ (less emphasis on B). (B) The density plot shows harms incurred by $4.8\times 10^{6}$ uniformly sampled designs in their response to archetype I (non-replicating, self-like/ auto-immune inducing or allergenic antigens; $R_{0}=1$) versus archetype II (immunogenic infection; high $\kappa$), as defined in Fig. 3B. Designs are sampled on a uniform Cartesian grid of sensitivities $\psi_{\sigma}\in[-3,3]$ and baseline rates $\ell_{i}\in[-3,3]$ (as in Fig. 3A). For many values of $\lambda$, we minimize the objective in (A), and trace out the Pareto front (magenta). For each point $\lambda$ on the front, we identify the ensemble of designs closest to the Pareto front ($\Theta(\lambda)$), where a design belongs to the near-Pareto ensemble at $\lambda$ if its incurred harm against archetypes I and II simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of the respective archetypes. The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-I harm incur larger increases in archetype-II harm (see SI Appendix, section 3). The average design parameter values over the near-elbow ensemble are $\langle\vec{\theta}\rangle_{\Theta(\lambda^{*})}=(2.89,1.67,-0.67,-0.22,-1.94,0.94)$. (C) Decomposition of total harm along the Pareto front in (B) into infection-related harm and toxicity (eq. 9). The black dashed line indicates the position of the elbow from (B). (D) Solid lines show parameter magnitudes for the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, as a function of harm incurred under archetype II along the Pareto curve (parameterized by $\lambda$). Shaded envelopes denote the standard deviation for the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front. (E) For each archetype in Fig. 3B, total harm is shown for the archetype-specific minimizer $\vec{\theta}^{(\mathrm{min})}$ (circle) and for the average over near-elbow ensemble in (B). See SI Appendix, Figs. S4, S5 for Pareto fronts formed by other pairs of archetypes. All simulation parameters other than the varied $K_{I}$, $b_{I}$, and $\vec{\theta}$ match Fig. 1.
Next, we explored regulatory design space by evaluating responses of different designs against diverse infection scenarios, sampled on a grid of immunogenicities ($\kappa$) and reproduction numbers $(R_{0})$ (Fig. 3B). For each infection scenario $(\kappa,R_{0})$, we identified the ensemble of near-optimal designs, $\Theta(\kappa,R_{0})$. A design $\vec{\theta}$ belongs to this ensemble if its incurred total harm (eq. 9) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection. Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, section 3). Four coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig. 3B): I—non-replicating, self-like/auto-immune inducing or allergenic antigens ($R_{0}=1$); II—highly immunogenic bacteria and viruses (high $\kappa$ or high $R_{0}$); III—intermediate replication and detectability, typifying mixed-strategy intracellular pathogens (mid-high $R_{0}$, mid-high $\kappa$); IV—difficult to detect threats such as growing tumors (low $\kappa$, broad range of $R_{0}$). The Pareto trade-off between infection harm and toxicity differs across these archetypes: designs optimized for archetypes I-III nearly eliminate infection damage with limited immunopathology, while designs in archetype IV show a strong trade-off (Fig. 3C).
Design parameters of effector programs minimizing harm vary across archetypes (Fig. 3D). For archetype I—non-replicating antigens—the optimal design is largely pre-programmed to suppress the immune response: signal sensitivities $\vec{\psi}$ are close to 0, and the baseline rates are tuned to slow activation, proliferation, and differentiation, and enforce rapid effector deletion. This damped response completely avoids toxicity when the threat does not grow so infection harm is determined by the initial exposure and cannot be decreased. Across all other archetypes a common pattern emerges—the response-derived cytokine signal $\sigma_{\text{res}}$ acts as an anti-inflammatory brake ($\psi_{\text{res}}<0$), with its magnitude tuned to context. Moreover, antigen sensitivity $\psi_{\text{Ag}}$ grows with immunogenicity: the effector program optimized for highly immunogenic pathogens in archetype II shows the largest $\psi_{\text{Ag}}$, whereas the programs for the poorly immunogenic archetype IV show diminished antigen sensitivity. Compared to archetype II, the baseline propensities $\ell$’s (or equivalently $g_{0}$’s) for archetype III accelerate effector contraction ($g_{0,E\to\emptyset}\,\uparrow$). For the weakly immunogenic archetype IV, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs. S2, S3).
Endogenous T-cell programs are shaped by minimizing harm from autoimmunity and acute infections
Although the relative evolutionary pressures exerted by different threats are unknown, we can still pose a principled multi-objective problem for the effector program, i.e., find a design $\vec{\theta}$ that minimizes the weighted sum of harms across pathogen archetypes:
$\vec{\theta}^{\prime}(\{\lambda_{i}\})=\underset{\vec{\theta}}{\arg\min}\sum_{i:\text{archetypes}}\lambda_{i}H^{(i)}(\vec{\theta}),$ (10)
where $\lambda_{i}\in[0,1]$ with the constraint $\sum_{i}\lambda_{i}=1$. The weights $\{\lambda_{i}\}$ determine the relative importance of each archetype in determining the optimal design $\vec{\theta}^{\prime}$. The solution to this family of problems forms a simplex over which $\lambda_{i}^{\prime}s$ vary (Fig. 4A, and SI Appendix, section 3). This approach accommodates uncertainty about the empirical prevalence of each type of infection and the preferences/contingencies of evolution.
Even though optimization can be done simultaneously over all four archetypes, we proceed by examining pairwise trade-offs that generate a one dimensional Pareto front, parameterized by $\lambda$ (Fig. 4A). A biologically salient case contrasts archetype I (non-replicating antigens) with archetype II (immunogenic pathogens). Non-replicating antigens plausibly arise from unprogrammed cell death events or external trauma inducing inflammation and may be implicated in autoimmune flares. This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Figs. S4, S5, for Pareto fronts formed by optimizing other pairs of clusters.
Fig. 4B shows harm from archetype I plotted against that of archetype II for a broad range of sampled designs and highlights the Pareto front (magenta). Designs close to the Pareto front achieve near-best total protection for a given compromise weight $\lambda$; moving to the right along the Pareto front shifts the emphasis from clearance of immunogenic pathogens ($\lambda\uparrow$) to tolerance ($\lambda\downarrow$). Decomposing total harm into infection-related harm and toxicity (eq. 9) shows that along this Pareto front it is relatively easy to reduce toxicity to non-replicating antigens without substantially worsening outcomes against immunogenic infections (Fig. 4C).
Figure 5: Elbow-optimal designs recapitulating macro-dynamics of T-cell expansion. (A,B) Fold expansion (i.e, peak clone size divided by the initial precursor number) of antigen-specific CD8+ T-cells in mice plotted against (A) the initial precursor number (data from [8]), and (B) the relative pMHC concentration required for half-maximal activation EC50 (data from [79]); in our model, EC50 is proxied by $K_{I}$. (C,D) Simulated mean fold expansion for designs belonging to the near-elbow ensemble from Fig. 4B ($\Theta(\lambda^{*})$) when their sensitivity to the infection-harm signal ($\psi_{\text{inf}}$) is varied (colors), shown as a function of (C) the initial precursor T-cell number, and (D) the immunogenicity of the antigen $\kappa$. For each design in the near-elbow ensemble identified in Fig. 4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$ and simulate responses against the broad sample of infections (Fig. 3B). Trend lines are obtained by averaging over infections and designs in the near-elbow ensemble (SI Appendix, Figs. S6, S7). Elbow designs with reduced sensitivity to infection-derived harm qualitatively match the experimental macro-dynamics of T-cell expansion. Unless otherwise indicated, simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}$.
Design Parameter Effect Source
baseline rates:
1. activation & prolif., $g_{N\to N^{*}},g_{E\to 2E}$ slow [72, 32]
2. differentiation, $g_{N^{*}\to E}$ slow [1]
3. contraction, $g_{E\to\emptyset}$ moderate [18, 32]
sensitivity to signal:
4. antigen, $\psi_{\text{Ag}}$ promote [72, 79, 49]
5. infection, $\psi_{\text{inf}}$ promote [10, 37, 54]
6. response, $\psi_{\text{res}}$ suppress [54, 17]
Table 1: Endogenous design of CD8+ T-cell response.
Next, we evaluated immune designs near the Pareto curve (Fig. 4D; SI Appendix, section 3). For each point $\lambda$, we identified an ensemble of designs near the Pareto front, $\Theta(\lambda)$. For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of those respective archetypes. The elbow of the Pareto curve represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype I will incur disproportionally large increases in harm from archetype II. The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$). Sensitivity to these signals diminishes as $\lambda$ decreases (i.e., as avoidance of autoimmunity is prioritized); see Fig. 4D. Baseline rates ($g_{0}$’s) of the designs at the elbow favor moderate to slow activation and proliferation of effectors, slow differentiation to effector state, and moderate contraction (Fig. 4D). As $\lambda$ increases (leftward; clearance weighed more), all baseline rates shift toward promoting more effector cells. As $\lambda$ decreases below the elbow (rightward), the already slow differentiation baseline rate falls toward zero, while activation/proliferation and contraction are tuned to suppress effector production, reflecting designs that minimize collateral damage even at the expense of incomplete clearance. Notably, despite being optimized only for archetypes I and II, designs in the near-elbow ensemble $\Theta(\lambda^{*})$ perform competitively for other archetypes, incurring harm close to the archetype-specific optimal designs (Fig. 4E; SI Appendix, Figs. S4, S5).
From these analyses alone, it is not possible to ascertain the compromises encoded in the endogenous T-cell response program. However, experiments on T-cell fate decisions in endogenous immunological programs (Table 1) are qualitatively consistent with designs that jointly minimize harm from immunogenic pathogens and auto-immune responses to transient injuries (elbow in Fig. 4B).
Figure 6: Engineering immune response for cancer immunotherapy. (A) The schematic shows the potential clearance of replicating cancer tumor cells by T-cells (e.g., infused during immunotherapy), and the potential collateral damage of therapy on susceptible tissue cells. The accessible levers in an immunotherapy include the dosage and antigenicity of the infused cells. Antigenicity is a property of the interaction between a specific TCR-antigen pair, and it can be increased by either changing the antigen or the TCR. Given a specific cognate T-cell, immunogenicity relates the antigenicity of a cancer cell to the antigenicity of a healthy cell. (B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm as a function of cancer antigenicity and dosage of administered T-cells. Yellow points correspond to the optimal dosage of T-cells for a given immunogenicity/antigenicity. (C) Total harm, harm from cancer and toxicity as functions of cancer immunogenicity ($\kappa$) are shown at the optimal T-cell dosage (yellow circles in B). (D) Clearance response—defined as the derivative of negative harm (i.e., protection) with respect to each design parameter $\theta_{i}$, $\partial H(\vec{\theta})/\partial{\theta_{i}}$ for $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$—averaged over the tuned near-elbow designs $\tilde{\Theta}(\lambda^{*})$, is shown for different levels of tumor immunogenicity $\kappa$; Left: $\vec{\psi}$ parameters; right: $\vec{g}_{0}$ parameters (colors). Shaded envelopes denote the standard deviation of outcomes over $\theta\in\tilde{\Theta}(\lambda^{*})$. For this analysis, we assume that T-cells are administered at high dosage ($N_{\text{lin.}}=10^{-2}S_{\text{max}}$). Simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.0\penalty 10000\ \text{day}^{-1},I_{0}=10^{-1}S_{\text{max}},b_{I}=0.1\penalty 10000\ \text{day}^{-1}$.
The choice of design impacts the dynamics of the immune response, which we can compare to experimental data. Prior experiments in mice [79, 8] indicate that peak fold-expansion of T-cell populations in response to a pathogen scales inversely with the initial number of cognate T-cells, and with the pMHC concentration for half-maximal response (proxied by $K_{I}$ in our model); see Fig. 5A, B. This pattern was recapitulated by explicitly modeling competition for antigens in polyclonal T-cell responses [53]. Near-elbow designs $\vec{\Theta}(\lambda^{*})$ (Fig. 4) reproduce this inverse dependence on initial clone size but underestimate the steepness of the trend with respect to immunogenicity ($\kappa$) (Fig. 5C, D). Reducing the sensitivity to infection $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamic range (Fig. 5D) with a minor impact on total harm compared to the effect of reducing sensitivity to antigen ($\psi_{\text{Ag}}$) for immunogenic infections (SI Appendix, Figs. S6, S7). This suggests that satisfying the observed macro-dynamic relationship is nearly orthogonal to the harm minimization objective. This effect cannot be achieved by tuning any other design parameter without relatively large increases in total harm (SI Appendix, Figs. S6, S7). However, we cannot rule out the possibility that sensitivity to infection harm signals is modulated by the strength of antigen signals. Such a mechanism might also explain the observed decline in expansion with declining immunogenicity.
To align responses with the observed macro-dynamic trends, subsequent analyses consider near-elbow designs $\vec{\theta}\in\Theta(\lambda^{*})$ with their infection-signal sensitivity reduced to one-tenth of their original value: for $\psi_{\text{inf}}\mapsto 0.1\,\psi_{\text{inf}}$. This ensemble of tuned designs is indicated by $\tilde{\Theta}(\lambda^{*})$. See SI Appendix, Figs. S9 and S10 for extensive model parameter sensitivity analyses with this ensemble.
Designing an effector response to a growing tumor
CD8+ T-cells are central to cancer surveillance and elimination of mutated pre-cancerous cells. We therefore asked how the tuned elbow-optimal designs $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$ identified for infectious threats (Fig. 4B, Fig. 5) respond to cancer, which we modeled as a growing, low-immunogenic tumor that kills healthy cells as it expands (Fig. 6A, SI Appendix, section 1).
Similar to the case of infection control, a clearance-toxicity trade-off arises, but it is sharpened in cancer because many tumor antigens are only marginally distinct from self. Even high-affinity recognition risks cross-reactivity and bystander damage. Tuned near-elbow designs incur high total harm for low-immunogenic tumors (Fig. 6B,C). As tumor immunogenicity increases, harm declines due to decreases in cross-reactive toxicity and increases in clearance (Fig. 6C).
Cancer immunotherapy leverages and improves upon the body’s own immune system to target tumors [9]. Patient-specific tumor-infiltrating lymphocytes (TILs) and TCR-engineered T-cells (TCR-T’s) purified and expanded ex-vivo, and infused in patients to recognize tumor peptide-MHC’s have shown clinical benefits. Chimeric antigen receptor (CAR) T-cells, on the other hand, target surface antigens without MHC restriction and can be deployed more broadly when targets are shared across patients. Contemporary CARs incorporate co-stimulatory domains (e.g., CD28 or 4-1BB) that enhance activation, proliferation, survival, and effector function [81]. In practice, three levers are pulled to engineer a T-cell therapy: (i) the antigen-recognition module (native TCR, TCR-T, or CAR) to specifically target the tumor; (ii) the intracellular signaling architecture (co-stimulation and engineered circuits) to sustains a robust and durable response; and (iii) the infused T-cell dose/schedule, tuned to maximize tumor control while limiting toxicity.
Our model predicts that the optimal dose should be tuned to tumor immunogenicity (Fig. 6B), consistent with prior findings on antigen-dependent T-cell dosing [62]. Examining the tuned near-elbow designs, we find: at low immunogenicity, clearance is inefficient and toxicity dominates, so smaller infusion doses minimize harm; at high immunogenicity, recognition and in-situ expansion are efficient, so even modest doses suffice; at intermediate immunogenicity, larger initial doses minimize total harm by accelerating tumor elimination while keeping toxicity tolerable (Fig. 6C).
TCR engineering raises an additional nuance: increasing TCR-tumor affinity often increases affinity to self [34], so that increasing tumor immunogenicity is difficult. To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged (SI Appendix, Fig. S8). We quantify affinity or reactivity of a cell by the ratio of the maximum healthy cell counts $S_{\text{max}}$ to the number of target cells required for half-maximal T-cell response (EC50), i.e., $K_{I}^{-1}S_{\text{max}}$ for tumor cells, and $K_{S}^{-1}S_{\text{max}}$ for self targets. Provided that tumor immunogenicity ($\kappa=K_{I}^{-1}K_{S}$) is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig. S8A) mirrors its dependence on immunogenicity in Fig. 6B. Nevertheless, these cross-reactive regimes (Fig. S8B) yield higher total harm than scenarios where self-interaction fixed so that increasing tumor antigenicity increases immunogenicity (Fig. 6C). This highlights the cost of affinity-driven cross-reactivity in TCR engineering for cancer therapy [45, 39, 73].
Finally, we asked which aspects of the endogenous CD8+ T-cell program could be modulated to improve tumor control. Using high-dose infusions for a broad range of immunogenicities, we computed the sensitivity of negative harm (i.e., protection) to small changes in design parameter values, i.e., $\nabla H(\vec{\theta})$ for $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$ (Fig. 6D). For very high-affinity T-cells (those able to respond at $\leq 1/100$ of the number of healthy cells), gains from further engineering are limited. For intermediate affinities—arguably more relevant clinically—several levers are impactful: increasing activation/proliferation baseline rates ($g_{0,N\to N^{*}},g_{0,E\to 2E}\uparrow$) [22] and decreasing effector death ($g_{0,E\to\emptyset}\downarrow$) provide strong improvements but may entail perturbing oncogene-linked pathways (e.g., MYC, BIM) and warrant caution [57]. Boosting antigen and infection sensitivities ($\psi_{\text{Ag}},\psi_{\text{inf}}\,\uparrow$), and tempering the anti-inflammatory brake ($|\psi_{\text{res}}|\downarrow$) also enhance clearance, albeit with a risk of runaway inflammation [69]. Increasing the baseline rate of differentiation ($g_{0,N^{*}\to E}\uparrow$) yields smaller gains. This suggests that reducing effector differentiation—e.g., via Tcf7 modulation [1, 22]—may provide a tractable knob to promote memory production and reduce relapse risk. Reduced effector differentiation may then be compensated for by tuning effector expansion and longevity.
Importantly, the achievable benefit of any perturbation for T-cell engineering will depend on patient- and tumor-specific constraints. These include the extent to which affinity-driven TCR cross-reactivity can be mitigated (Fig. S8) and may require patient-specific measurements of T-cell interactions and signaling to calibrate designs for robust, durable therapy.
Discussion
In this work, we modeled the CD8+ T-cell response as a feedback-controlled program that integrates three cues—instantaneous antigen, cumulative infection-derived harm, and cumulative response-derived harm—to regulate activation, expansion, differentiation, and contraction of T-cells. Modulation of response proportional to the antigen signal and the time-integrated harm signals mirrors proportional-integral-derivative (PID) control of engineered systems, a strategy known for its stable and robust steering of dynamical processes toward desired outcomes [7].
Exploring an ensemble of biologically motivated controllers (designs) and optimizing for pathogen clearance while minimizing immunopathology exposed a Pareto trade-off: a reduction in infection harm generally leads to an increased collateral damage to healthy tissues. This pattern persisted across diverse pathogenic challenges defined by their immunogenicity and within-host basic reproduction number.
Three design principles emerged: (i) Pre-programmed (baseline-tuned) designs help restrain autoimmunity and slowly proliferating threats but are inefficient against fast-proliferating infections or metastatic tumors; effective control there requires signal integration and feedback. (ii) Among signal sources, antigen and response-harm dominate; sensitivity to infection-harm contributes comparatively little once CD8+ effector T-cell expansion is underway. (iii) Feedback from response-harm acts as an anti-inflammatory brake, enabling strong antigen-driven responses while limiting collateral damage when the response overshoots—a characteristic consistent with endogenous effector programs.
We posited that effector programs should attempt to solve a multi-objective problem to counter diverse pathogenic challenges. Interestingly, we found that a Pareto-optimal design that suppresses responses to frequent, self-like injuries while clearing recognizable, replicating (e.g., acute) infections mirrors the endogenous T-cell programs in vertebrates: strongly antigen-driven with an anti-inflammatory brake, and conservative cell-fate transition speeds that prevent runaway effector responses.
We assumed that, although memory formation is essential, its immediate impact on primary effector magnitude is modest: $5-10\%$ of peak effector numbers [38, 1]. Memory “quality” (longevity, proliferative capacity) depends on differentiation timing and the avidity of the selected clones [12, 78, 40, 27]. To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall. An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the hosts’ evolutionary survival pressures that could have shaped the effector program.
Using a minimal tumor model, we showed how the design objective informs rational immunotherapy protocols. First, it suggests that infused T-cell dosage should depend on tumor immunogenicity—balancing clearance against collateral damage, as previously emphasized in the literature [9, 39, 62]. Second, treating the design parameters as engineering targets yields testable strategies to improve efficacy and safety, which is highly desirable [76, 44, 5]. For moderately immunogenic tumors, promoting effector output by raising baseline rates of activation/proliferation or increasing antigen and infection sensitivities, while tempering the anti-inflammatory brake, can produce robust responses.
To operationalize these ideas, abstract signal sensitivities and baseline rates must be mapped to gene programs as targets of cell engineering (Table 1). Combining principled control-theoretic design with patient-specific measurements of antigenicity and damage signaling may enable tailored effector programs that respect the fundamental trade-offs of response from this study, while bending them toward better outcomes [47].
Key questions follow from our study. Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs? Endogenous immune responses are polyclonal [68]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses? Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them? Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
II Data and code availability
Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune.
Acknowledgment
This work has been supported by the CAREER award from the National Science Foundation grant 2045054, and the National Institutes of Health MIRA award R35 GM142795. This work is also supported, in part, through the Departments of Physics and Applied Mathematics and the College of Arts and Sciences at the University of Washington. The numerical analyses in this work were completed on Hyak, the UW’s high performance computing cluster, which is funded by the UW student technology fee. This work benefited from discussions during the 2024 program “Interactions and Co-evolution between Viruses and Immune Systems” at the Kavli Institute for theoretical physics (KITP), which is supported by National Science Foundation grant PHY-2309135, and the Gordon and Betty Moore Foundation Grant No. 2919.02.

### v2

@@
 Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive.
-Here, we recast a T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues.
-We demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings.
+Here, we recast the T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues.
+By exploring a broad class of feedback-controller designs as potential immune programs, we demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings.
 We recover an inherent trade-off: infection clearance at the cost of immunopathology.
 We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals.
-This model explains experimentally observed patterns of T-cell effector expansion regulation in mice.
+Notably, we find that designs that balance harm from acute infections and autoimmunity produce immune responses consistent with experimentally observed patterns of T-cell effector expansion in mice.
 Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment.
@@
 We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology.
-Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [28, 41, 17, 54], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses.
+Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [25, 26, 27, 28, 29], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses.
 Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms.
@@
 Formulating and testing principled heuristics provides a baseline to gauge the performance of real immune systems in responding to infections.
-This perspective has yielded deep insights in other biological contexts, revealing the architectural logic of genetic and regulatory networks [23, 67, 42, 31, 6].
-In immunology, high-level mathematical models have demonstrated how a finite lymphocyte repertoire can be distributed to best confront a diverse pathogenic landscape [59, 50, 11] and how immune memory should be allocated to counter future threats [51, 64, 15, 65].
+This perspective has yielded deep insights in other biological contexts, revealing the architectural logic of genetic and regulatory networks [30, 31, 32, 33, 34, 35].
+In immunology, high-level mathematical models have demonstrated how a finite lymphocyte repertoire can be distributed to best confront a diverse pathogenic landscape [36, 37, 38] and how immune memory should be allocated to counter future threats [39, 40, 41, 42, 43].
 By extending this perspective to the dynamics of CD8+ T-cell signal integration and response, our work provides a quantitative framework that links observed decision-making to operating principles and trade-offs.
@@
 Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$.
-In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates.
+In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates (Fig.
+S1).
 Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory.
@@
 We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells.
-Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Fig.
-1A).
+Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs.
+1A, S1).
 The dynamics of the susceptible and infected cells follow,
 $\begin{split}\dot{S}&=-b_{I}IS-d_{E}\frac{E}{K_{S}+I\frac{K_{S}+E}{K_{I}+E}+E+S}S\\ \dot{I}&=b_{I}IS-d_{I}I-d_{E}\frac{E}{K_{I}+S\frac{K_{I}+E}{K_{S}+E}+E+I}I,\end{split}$ (1)
-where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are inverse affinity constants quantifying how well effector cells bind to and kill infected and susceptible cells, respectively [14, 25, 30].
-The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix, section 1).
+where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [44, 45, 46].
+The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix).
 We quantify within-host reproductive success of a pathogen by the intra-host basic reproduction number,
@@
 We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells.
-These inverse affinity parameters determine the immunogenicity of an infection, which we define as,
+These recognition thresholds determine the immunogenicity of an infection, which we define as,
 $\text{immunogenicity:}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \kappa=K_{I}^{-1}K_{S}.$ (3)
 This dimensionless quantity reflects how readily cognate T-cells recognize antigens presented by infected cells as “non-self” relative to those presented by healthy cells (self).
-We consider infection scenarios with broad ranges of immunogenicity ($\kappa$) and basic reproduction numbers ($R_{0}$) (SI Appendix, section 1).
+We consider infection scenarios with broad ranges of immunogenicity ($\kappa$) and basic reproduction numbers ($R_{0}$) (SI Appendix).
 Signal cues to modulate T-cell decision-making.
@@
 Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$.
-Together, expressed by their inverse affinity constants, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
+Together, expressed by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
 $\begin{split}&\text{antigen signal:}\penalty 10000\ \penalty 10000\ \sigma_{\text{Ag}}(t)={K_{I}}^{-1}{I}(t)\\ &\text{harm signals:}\penalty 10000\ \penalty 10000\ \sigma_{\text{inf}/\text{res}}(t)=K_{H}^{-1}\int_{0}^{t}e^{-(t-s)d_{H}}\,h_{\text{inf}/\text{res}}(s)\,{\rm d}s.\end{split}$ (5)
-The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix, section 1).
+The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).
 By disentangling these sources of (potentially redundant) signals, we probe how CD8+ T-cells use these cues to make fate decisions.
@@
 Naïve T-cells cycle through lymphoid tissues, becoming activated when they durably engage cognate peptide-MHC complexes on activated APCs [49, 50].
-Activated T-cells can differentiate into an effector phenotype, which participates in infection clearance, or into a long-lived memory state (Fig.
-1A).
-Here, we examine the course of CD8+ T-cells response during an infection, dividing the process into four canonical response modules (Figure.
+Activated T-cells can differentiate into an effector phenotype, which participates in infection clearance, or into a long-lived memory state (Figs.
+1A, S1).
+Here, we examine the course of CD8+ T-cells response during an infection, dividing the process into four canonical response modules (Fig.
 1A):
@@
 $\begin{split}&g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi}_{i},\ell_{i})=\frac{1}{1+\exp\left[-L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})\right]},\\ &\text{with,}\\ &L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})=\ell_{i}+\sum_{\alpha\in\{\text{Ag},\text{inf},\text{res}\}}\psi^{\alpha}_{i}\log\left[1+{\sigma_{\alpha}}\right].\end{split}$ (7)
-The parameter $\ell_{i}$ sets the baseline propensity of transition $i$: in the absence of external signals, the regulatory factor evaluates to $g_{0,i}=(1+\exp(-\ell_{i}))^{-1}$ (Figure.
+The parameter $\ell_{i}$ sets the baseline propensity of transition $i$: in the absence of external signals, the regulatory factor evaluates to $g_{0,i}=(1+\exp(-\ell_{i}))^{-1}$ (Fig.
 1C).
@@
 Figure 2: Design parameters shape T-cell response and infection clearance.
-(A) Varying the design parameters influences effector trajectories, summarized by four features: activation time, expansion rate, peak effector magnitude, and contraction rate.
-(B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom)—measured in units of maximum number of susceptible cells $S_{\text{max}}$—for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel.
+(A) Varying the design parameters influences effector trajectories.
+(B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom), measured in units of maximum number of susceptible cells $S_{\text{max}}$, for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel.
 All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$.
 Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig.
-S1.
-(C) Sensitivities of trajectory (as in B) fold expansion, defined as the maximum effector count divided by the initial number of T-cells $N_{\text{lin.}}$, to changes in baseline transition rates (colors).
+S2.
+(C) Peak effector count, measured in units of maximum number of susceptible cells $S_{\text{max}}$, is shown by varying each of the three baseline transition rates (colors), while keeping the rest of the parameters fixed (similar to (B)).
 Parameters are the same as Fig.
@@
 We systematically sample $\vec{\theta}$’s to assess performance of designs against different infections (Fig.
-1D; SI Appendix, section 3).
-While $\vec{\theta}$ is broadly sampled, we hold the remaining biological rates and network parameters underlying the decision-making program in Fig.
+1D; SI Appendix).
+While $\vec{\theta}$ is broadly sampled, we hold the remaining biological rates and design parameters underlying the decision-making program in Fig.
 1A within empirically supported ranges reported in SI Appendix, Table S1.
@@
 $\begin{split}H(\vec{\theta})=\underbrace{\int_{0}^{T}\left[h_{\text{inf}}\left(t,\vec{\theta}\right)+h_{E,I}\left(t,\vec{\theta}\right)\right]{\rm d}t}_{H_{\rm inf.}(\vec{\theta})}+\underbrace{\int_{0}^{T}h_{E,S}\left(t,\vec{\theta}\right){\rm d}t}_{H_{\rm toxic.}(\vec{\theta})},\end{split}$ (9)
-where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, section 1 and 2, and Table S1).
+where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, Table S1).
 This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig.
@@
 (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A).
-By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I, II, III, IV (SI Appendix, section 3).
+By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I–IV, as indicated (SI Appendix).
 (C) Toxicity vs.
@@
 Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated cytokine signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig.
-S1).
+S2).
 Effector designs trade infection harm for toxicity
 Previous research has identified a trade-off between maximizing infection clearance and minimizing response toxicity (immunopathology) in the immune response [61, 28, 29].
-Consistently, by systematically sampling the space of designs (SI Appendix, section 3), we found this trade-off in the form of a Pareto front: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig.
+Consistently, by systematically sampling the space of designs (SI Appendix), we found this trade-off in the form of a Pareto front: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig.
 3A).
@@
 This trade-off is stronger for infections with a higher intra-host reproductive number $R_{0}$ or lower immunogenicity $\kappa$.
-Interestingly, for scenarios with lower $R_{0}$, many of the hard-wired networks without signal feedback ($\vec{\psi}_{i}=0$) lie along the Pareto front, and perform near-optimally.
+Interestingly, for scenarios with lower $R_{0}$, many of the hard-wired response programs without signal feedback ($\vec{\psi}_{i}=0$) lie along the Pareto front, and perform near-optimally.
 For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far inside the sub-optimal region.
@@
 For each point $\lambda$ on the front, we identify the ensemble of designs closest to the Pareto front ($\Theta(\lambda)$), where a design belongs to the near-Pareto ensemble at $\lambda$ if its incurred harm against archetypes I and II simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of the respective archetypes.
-The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-I harm incur larger increases in archetype-II harm (see SI Appendix, section 3).
+The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-I harm incur larger increases in archetype-II harm (see SI Appendix).
 The average design parameter values over the near-elbow ensemble are $\langle\vec{\theta}\rangle_{\Theta(\lambda^{*})}=(2.89,1.67,-0.67,-0.22,-1.94,0.94)$.
@@
 (D) Solid lines show parameter magnitudes for the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, as a function of harm incurred under archetype II along the Pareto curve (parameterized by $\lambda$).
-Shaded envelopes denote the standard deviation for the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front.
+Shaded envelopes denote the standard deviation of the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front.
 (E) For each archetype in Fig.
@@
 See SI Appendix, Figs.
-S4, S5 for Pareto fronts formed by other pairs of archetypes.
+S6, S7 for Pareto fronts formed by other pairs of archetypes.
 All simulation parameters other than the varied $K_{I}$, $b_{I}$, and $\vec{\theta}$ match Fig.
@@
 9) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection.
-Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, section 3).
+Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, Fig.
+S3).
 Four coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig.
@@
 For the weakly immunogenic archetype IV, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs.
-S2, S3).
+S4, S5).
 Endogenous T-cell programs are shaped by minimizing harm from autoimmunity and acute infections
@@
 The solution to this family of problems forms a simplex over which $\lambda_{i}^{\prime}s$ vary (Fig.
-4A, and SI Appendix, section 3).
+4A; SI Appendix).
 This approach accommodates uncertainty about the empirical prevalence of each type of infection and the preferences/contingencies of evolution.
@@
 This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Figs.
-S4, S5, for Pareto fronts formed by optimizing other pairs of clusters.
+S6, S7, for Pareto fronts formed by optimizing other pairs of clusters.
 Fig.
@@
 For each design in the near-elbow ensemble identified in Fig.
-4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$ and simulate responses against the broad sample of infections (Fig.
+4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$ (colors), and simulate responses against the broad sample of infections (Fig.
 3B).
 Trend lines are obtained by averaging over infections and designs in the near-elbow ensemble (SI Appendix, Figs.
-S6, S7).
+S8, S9).
 Elbow designs with reduced sensitivity to infection-derived harm qualitatively match the experimental macro-dynamics of T-cell expansion.
 Unless otherwise indicated, simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}$.
+Table 1: Endogenous design of CD8+ T-cell response.
 Design Parameter Effect Source
 baseline rates:
-1. activation & prolif., $g_{N\to N^{*}},g_{E\to 2E}$ slow [72, 32]
-2. differentiation, $g_{N^{*}\to E}$ slow [1]
+1. activation & prolif., $g_{N\to N^{*}},g_{E\to 2E}$ low [4, 17]
+2. differentiation, $g_{N^{*}\to E}$ low [16]
 3. contraction, $g_{E\to\emptyset}$ moderate [63, 17]
@@
 6. response, $\psi_{\text{res}}$ suppress [29, 28]
-Table 1: Endogenous design of CD8+ T-cell response.
 Next, we evaluated immune designs near the Pareto curve (Fig.
-4D; SI Appendix, section 3).
+4D; SI Appendix).
 For each point $\lambda$, we identified an ensemble of designs near the Pareto front, $\Theta(\lambda)$.
 For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of those respective archetypes.
-The elbow of the Pareto curve represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype I will incur disproportionally large increases in harm from archetype II.
+The elbow of the Pareto curve represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype I will incur disproportionally large increases in harm from archetype II (see SI Appendix).
 The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$).
@@
 4E; SI Appendix, Figs.
-S4, S5).
+S6, S7).
 From these analyses alone, it is not possible to ascertain the compromises encoded in the endogenous T-cell response program.
@@
 Antigenicity is a property of the interaction between a specific TCR-antigen pair, and it can be increased by either changing the antigen or the TCR.
-Given a specific cognate T-cell, immunogenicity relates the antigenicity of a cancer cell to the antigenicity of a healthy cell.
+Given a specific cognate T-cell, immunogenicity is the ratio of the antigenicity of a cancer cell to that of a healthy cell.
 (B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm as a function of cancer antigenicity and dosage of administered T-cells.
@@
 5C, D).
-Reducing the sensitivity to infection $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamic range (Fig.
-5D) with a minor impact on total harm compared to the effect of reducing sensitivity to antigen ($\psi_{\text{Ag}}$) for immunogenic infections (SI Appendix, Figs.
-S6, S7).
-This suggests that satisfying the observed macro-dynamic relationship is nearly orthogonal to the harm minimization objective.
-This effect cannot be achieved by tuning any other design parameter without relatively large increases in total harm (SI Appendix, Figs.
-S6, S7).
-However, we cannot rule out the possibility that sensitivity to infection harm signals is modulated by the strength of antigen signals.
-Such a mechanism might also explain the observed decline in expansion with declining immunogenicity.
+Reducing the sensitivity to infection harm $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamic range (Fig.
+5D) while having only a minor effect on total harm, and hence, our optimization objective (Fig.
+S9).
+Interestingly, adjusting the other sensitivities ($\psi_{\text{Ag}},\psi_{\text{res}}$), or the baseline rates $g_{0}$’s does not reproduce the dynamic range observed in the experiments (Figs.
+S8, S9).
 To align responses with the observed macro-dynamic trends, subsequent analyses consider near-elbow designs $\vec{\theta}\in\Theta(\lambda^{*})$ with their infection-signal sensitivity reduced to one-tenth of their original value: for $\psi_{\text{inf}}\mapsto 0.1\,\psi_{\text{inf}}$.
@@
 See SI Appendix, Figs.
-S9 and S10 for extensive model parameter sensitivity analyses with this ensemble.
+S11 and S12 for extensive model parameter sensitivity analyses with this ensemble.
 Designing an effector response to a growing tumor
@@
 5) respond to cancer, which we modeled as a growing, low-immunogenic tumor that kills healthy cells as it expands (Fig.
-6A, SI Appendix, section 1).
+6A; SI Appendix).
 Similar to the case of infection control, a clearance-toxicity trade-off arises, but it is sharpened in cancer because many tumor antigens are only marginally distinct from self.
@@
 To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged (SI Appendix, Fig.
-S8).
+S10).
 We quantify affinity or reactivity of a cell by the ratio of the maximum healthy cell counts $S_{\text{max}}$ to the number of target cells required for half-maximal T-cell response (EC50), i.e., $K_{I}^{-1}S_{\text{max}}$ for tumor cells, and $K_{S}^{-1}S_{\text{max}}$ for self targets.
 Provided that tumor immunogenicity ($\kappa=K_{I}^{-1}K_{S}$) is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig.
-S8A) mirrors its dependence on immunogenicity in Fig.
+S10A) mirrors its dependence on immunogenicity in Fig.
 6B.
 Nevertheless, these cross-reactive regimes (Fig.
-S8B) yield higher total harm than scenarios where self-interaction fixed so that increasing tumor antigenicity increases immunogenicity (Fig.
+S10B) yield higher total harm than scenarios where self-interaction fixed so that increasing tumor antigenicity increases immunogenicity (Fig.
 6C).
@@
 Reduced effector differentiation may then be compensated for by tuning effector expansion and longevity.
-Importantly, the achievable benefit of any perturbation for T-cell engineering will depend on patient- and tumor-specific constraints.
+Interestingly, recent CRISPR screens show that among other candidates, knocking out FAS and PTPN2 genes in CAR T-cells reduces apoptosis ($g_{0,E\to\emptyset}\downarrow$) and redirects differentiation toward effector programs ($g_{0,E\to 2E}\uparrow$), thereby enhancing CAR T-cell efficacy [76, 77, 78, 79].
+Our modeling framework can help prioritize targets for such screens to rationally identify candidate genes that improve T-cell immunotherapies.
+It should be noted that the achievable benefit of any perturbation for T-cell engineering will depend on patient- and tumor-specific constraints.
 These include the extent to which affinity-driven TCR cross-reactivity can be mitigated (Fig.
-S8) and may require patient-specific measurements of T-cell interactions and signaling to calibrate designs for robust, durable therapy.
+S10) and may require patient-specific measurements of T-cell interactions and signaling to calibrate designs for robust, durable therapy.
 Discussion
@@
 To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall.
-An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the hosts’ evolutionary survival pressures that could have shaped the effector program.
+An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the host’s evolutionary survival pressures that could have shaped the effector program.
 Using a minimal tumor model, we showed how the design objective informs rational immunotherapy protocols.
@@
 Key questions follow from our study.
-Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs?
+Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs [88, 89]?
 Endogenous immune responses are polyclonal [90]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses?
@@
 Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune.
-Acknowledgment
-This work has been supported by the CAREER award from the National Science Foundation grant 2045054, and the National Institutes of Health MIRA award R35 GM142795.
+Acknowledgement
+This work has been supported by the CAREER award from the National Science Foundation grant 2045054 (A.N., O.A.U., Z.M.), the National Institutes of Health MIRA award R35 GM142795 (A.N.), and the Intramural Research Program of the National Cancer Institute, Project # ZIA BC 012007 (G.A.-B.).
 This work is also supported, in part, through the Departments of Physics and Applied Mathematics and the College of Arts and Sciences at the University of Washington.
 The numerical analyses in this work were completed on Hyak, the UW’s high performance computing cluster, which is funded by the UW student technology fee.
-This work benefited from discussions during the 2024 program “Interactions and Co-evolution between Viruses and Immune Systems” at the Kavli Institute for theoretical physics (KITP), which is supported by National Science Foundation grant PHY-2309135, and the Gordon and Betty Moore Foundation Grant No.
+This work benefited from discussions during the 2023 summer workshop “Statistical Physics and Adaptive Immunity” at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and the 2024 program “Interactions and Co-evolution between Viruses and Immune Systems” at the Kavli Institute for theoretical physics (KITP), which is supported by National Science Foundation grant PHY-2309135, and the Gordon and Betty Moore Foundation Grant No.
 2919.02.

### v3

@@
 Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive.
-Here, we recast the T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues.
+Here, we recast T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues.
 By exploring a broad class of feedback-controller designs as potential immune programs, we demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings.
@@
 We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals.
-Notably, we find that designs that balance harm from acute infections and autoimmunity produce immune responses consistent with experimentally observed patterns of T-cell effector expansion in mice.
-Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment.
+Notably, we find the designs that balance the harm from acute infections and autoimmunity produce immune responses consistent with the experimentally observed patterns of T-cell effector expansion in mice.
+Extending our model to immune-based T-cell therapies for cancer tumors, we quantify the trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment.
 Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells.
@@
 Specifically, we cast each CD8+ T-cell as a feedback controller whose intra-cellular regulatory program modulates activation, differentiation, proliferation, and death in response to intra-host cues, such as antigen levels and cytokine signals, which themselves are dynamic in the course of an infection [24, 7].
-We then ask how alternative immunological objectives give rise to distinct effector programs.
+We then ask how alternative immunological contexts require distinct effector programs.
 First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules.
@@
 Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms.
-Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments—and thus candidate genetic engineering—that could enhance T-cell-based cancer immunotherapy.
+Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments through genetic engineering that could enhance T-cell-based cancer immunotherapy.
 The immune system’s true performance objective, and how evolution has fine-tuned it, may ultimately be unknowable.
@@
 Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$).
-Infection dynamics $(S,I)$ are modeled by ODEs coupled to a stochastic dynamics of the T-cell response.
+Infection dynamics $(S,I)$ are modeled by ODEs coupled to stochastic dynamics of the T-cell response.
 Memory transitions (gray arrows) are neglected in our model.
 (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time.
+Cell counts are reported in units of the maximum number of susceptible cells $S_{\text{max}}$.
 (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs.
-6, 7).
+4, 5).
 The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq.
-8).
+6).
 (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq.
-2) and immunogenicities $\kappa$ (eq.
-3).
+1) and immunogenicities $\kappa$ (eq.
+2).
 For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm).
 Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq.
-9).
-Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
-Model
+7).
+Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},T=30\penalty 10000\ \text{days},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
+II Model
 To characterize the immune response to an infection, we model the large-population dynamics of susceptible and infected cells in a host as deterministic processes, coupled to a stochastic, time-inhomogeneous birth-death process for the CD8+ T-cell response (Fig.
-1A, and SI Appendix for details).
+1A, B and SI Appendix).
 Intra-host infection dynamics.
-We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells.
-Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs.
-1A, S1).
-The dynamics of the susceptible and infected cells follow,
-$\begin{split}\dot{S}&=-b_{I}IS-d_{E}\frac{E}{K_{S}+I\frac{K_{S}+E}{K_{I}+E}+E+S}S\\ \dot{I}&=b_{I}IS-d_{I}I-d_{E}\frac{E}{K_{I}+S\frac{K_{I}+E}{K_{S}+E}+E+I}I,\end{split}$ (1)
-where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [44, 45, 46].
-The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix).
+We model a population of healthy, susceptible cells $S(t)$.
+Pathogen exposure seeds an infected subpopulation $I(t)$, which grows over time $t$ via encounters with an otherwise homeostatically stable susceptible cell population, and declines through infected-cell death, triggering an immune response (Figs.
+1A, B S1).
+Activated effector T-cells $E(t)$ can kill both infected and susceptible cells, with clearance governed by recognition thresholds $K_{I}$ and $K_{S}$, respectively, which quantify how readily effector cells recognize each cellular subpopulation; higher thresholds indicate that more effector cells are needed for recognition and killing.
+We model this dynamics as a deterministic birth-death process with competitive interaction of T-cells with infected and healthy cells; see Materials and Methods and the SI Appendix.
 We quantify within-host reproductive success of a pathogen by the intra-host basic reproduction number,
 $R_{0}=\frac{b_{I}S_{\text{max}}}{d_{I}},$ (1)
-measuring the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population.
+where $b_{I}/d_{I}$ is the ratio of birth to death rates for infected cells, and $S_{\text{max}}$ is the maximum number of susceptible cells in a tissue (Materials and Methods).
+$R_{0}$ measures the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population.
 This definition mirrors the standard epidemiological reproduction number [44, 45].
-We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells.
-These recognition thresholds determine the immunogenicity of an infection, which we define as,
+We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume that the T-cell recognition threshold for infected cells is lower than that for susceptible cells ($K_{I}\leq K_{S}$), such that T-cells more readily respond to infected cells.
+Accordingly, we define the immunogenicity of an infection as,
 $\text{immunogenicity:}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \kappa=K_{I}^{-1}K_{S}.$ (2)
@@
 $\begin{split}h(t)&=h_{\text{inf}}(t)+\underbrace{h_{E,I}(t)+h_{E,S}(t)}_{=h_{\text{res}}(t)},\end{split}$ (3)
-where $h_{\text{inf}}(t)$ represents the number of cells killed directly by the pathogen (infection), and $h_{E,I}(t)$ and $h_{E,S}(t)$ denote the number of infected and healthy cells, respectively, killed by the effector T-cell response, amounting to the total damage by the response $h_{\text{res}}(t)$.
+where $h_{\text{inf}}(t)$ represents the number of cells killed directly by the pathogen (infection), and $h_{E,I}(t)$ and $h_{E,S}(t)$ denote the number of infected and healthy cells, respectively, killed by the effector T-cell response, amounting to the total damage from the response $h_{\text{res}}(t)$.
 […]
 We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively.
-Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$.
-Together, expressed by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
-$\begin{split}&\text{antigen signal:}\penalty 10000\ \penalty 10000\ \sigma_{\text{Ag}}(t)={K_{I}}^{-1}{I}(t)\\ &\text{harm signals:}\penalty 10000\ \penalty 10000\ \sigma_{\text{inf}/\text{res}}(t)=K_{H}^{-1}\int_{0}^{t}e^{-(t-s)d_{H}}\,h_{\text{inf}/\text{res}}(s)\,{\rm d}s.\end{split}$ (5)
-The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).
+Presented antigens on APCs serve as an additional cue.
+We take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$, which together with the other two signals, modulate the cellular decision-making processes in Fig.
+1A; see Materials and Methods and SI Appendix for details.
 By disentangling these sources of (potentially redundant) signals, we probe how CD8+ T-cells use these cues to make fate decisions.
-We do not pre-assign any of the signals to a pro- or anti-inflammatory role but allow optimization goals to determine their roles across diverse infection scenarios.
+We do not pre-assign any signals to a pro- or anti-inflammatory role but allow optimization goals to determine their influence across diverse infection scenarios.
+We should emphasize that CD8+ responses do not occur in isolation: the innate immune system, regulatory T-cells (Tregs), helper T-cells, and B-cells can all directly or indirectly shape CD8+ dynamics.
+While these interactions are not modeled explicitly, we interpret the harm and antigen signals as coarse-grained proxies for inputs from other immune compartments, capturing their net pro- and anti-inflammatory effects.
+In the same vein, our immunogenicity parameter $\kappa$ (eq.
+2) may also subsume regulatory constraints, e.g.
+through Treg-mediated suppression that limits effector responses to self-like antigens, which in our framework correspond to low-immunogenicity targets.
+Incorporating these additional components explicitly would be an important direction for future work, enabling the study of richer feedbacks and more granular immune regulation and dynamics.
 Design of T-cell programs for infection response.
@@
     Contraction: programmed loss of effector cells, $E\to\emptyset$.
-These response processes map onto partially distinct genetic programs: the MYC pathway drives activation and proliferation [17]; the Wnt-Tcf7 pathway regulates effector versus memory fate decisions [51, 16]; and the Bcl-2-family of proteins control contraction of the effector cells via apoptosis [17].
+These response processes map onto partially distinct genetic programs: the MYC pathway drives activation and proliferation [17]; the Wnt-Tcf7 pathway regulates effector versus memory fate decisions [48, 16]; contraction can be attributed to effector cell apoptosis controlled by the Bcl-2-family of proteins, or the differentiation of effectors into effector-memory cells [16].
+The latter is regulated by the balance of T-bet/Blimp-1 (effector) versus Bcl-6/Eomes (memory) programs, as well as metabolic pathways such as mTORC1 [49].
+To reduce model degeneracy, we do not distinguish among these distinct biological sources of contraction in our analysis.
 For each process $i$, we represent its genetic regulation by a time-dependent transition rate $r_{i}(t)$, modulated by the instantaneous signals $\vec{\sigma}(t)$ (eq.
-5) received by the T-cell (Fig.
+10) received by the T-cell (Fig.
 1C),
 $r_{i}(t)=$ $r^{\text{max}}\ g_{i}\left(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell_{i}\right),$ (4)
-where the maximal rate $r^{\text{max}}$ is capped by the cell-cycle time [13, 21, 49], and $g_{i}$ is a monotonic response function of signal input, chosen from the modified Monod-Wyman-Changeux functions (Hill-type family) [52, 53]:
+where the maximal rate $r^{\text{max}}$ is capped by the cell-cycle speed [13, 21, 46], and $g_{i}$ is a monotonic response function of signal input, chosen from the modified Monod-Wyman-Changeux functions (Hill-type family) [50, 51]:
 $\begin{split}&g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi}_{i},\ell_{i})=\frac{1}{1+\exp\left[-L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})\right]},\\ &\text{with,}\\ &L_{i}(\vec{\sigma};\vec{\psi}_{i},\ell_{i})=\ell_{i}+\sum_{\alpha\in\{\text{Ag},\text{inf},\text{res}\}}\psi^{\alpha}_{i}\log\left[1+{\sigma_{\alpha}}\right].\end{split}$ (5)
@@
 (A) Varying the design parameters influences effector trajectories.
-(B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom), measured in units of maximum number of susceptible cells $S_{\text{max}}$, for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel.
+(B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom) for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel.
+Both quantities are measured in units of the maximum number of susceptible cells, $S_{\text{max}}$, which is a convenient reference scale for different cell sub-populations.
 All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$.
@@
 (C) Peak effector count, measured in units of maximum number of susceptible cells $S_{\text{max}}$, is shown by varying each of the three baseline transition rates (colors), while keeping the rest of the parameters fixed (similar to (B)).
+Solid lines show the mean of effector counts across $n=5$ simulation replicates and shaded envelopes indicate one standard deviation (Materials and Methods).
 Parameters are the same as Fig.
@@
 We therefore adopt a cell-economics objective in which harm is measured by the total number of cells lost over the course of infection.
-We evaluate the performance of a given design $\vec{\theta}$ by simulating the response dynamics (Fig. 1D) and computing the total cell death (or inflicted harm) as
+We evaluate the performance of a given design $\vec{\theta}$ by simulating the response dynamics (Fig.
+1D) and computing the total cell death (or inflicted harm) as
 $\begin{split}H(\vec{\theta})=\underbrace{\int_{0}^{T}\left[h_{\text{inf}}\left(t,\vec{\theta}\right)+h_{E,I}\left(t,\vec{\theta}\right)\right]{\rm d}t}_{H_{\rm inf.}(\vec{\theta})}+\underbrace{\int_{0}^{T}h_{E,S}\left(t,\vec{\theta}\right){\rm d}t}_{H_{\rm toxic.}(\vec{\theta})},\end{split}$ (7)
-where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, Table S1).
-This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig.
+where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the fixed time window $T$.
+In our main simulations, we set $T=30$ days to reflect the typical timescale of acute infections and the associated immune response (SI Appendix, Table S1), but we also assess the sensitivity of our results to variation in $T$ (Materials and Methods).
+This single, interpretable metric of harm (eq.
+7) balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig.
 1D).
@@
 (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq.
-9)—measured in units of $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)).
+7)—measured in units of the maximum number of susceptible cells $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)).
 Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq.
-8).
+6).
 Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed.
@@
 (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A).
-By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I–IV, as indicated (SI Appendix).
+By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify three pathogen “archetypes” (colors): I–III, as indicated (SI Appendix).
 (C) Toxicity vs.
-infection harm associated with optimal designs for each archetype is shown.
-Large dots represent cluster averages and small transparent dots represent individual infections, sampled in (B).
-(D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for archetype $i$ to within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)).
+infection harm associated with the optimal harm-minimizing designs for each pathogen archetype in (B) is shown.
+Large points indicate archetype means; small semi-transparent points show individual immune challenges sampled in (B).
+Reducing infection harm (horizontal axis) comes at the expense of increased toxicity (vertical axis).
+(D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for each archetype within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)).
 Error bars denote the standard deviation within this ensemble.
 All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$).
-Unless otherwise noted, simulation parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$
-Results
+Unless otherwise noted, results represent the average of $n=5$ simulation replicates with parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$
+III Results
 Cell-fate transitions determine effector response and pathogen clearance
 How do rates of cell-fate transitions, determined by the regulatory design $\vec{\theta}$ (eq.
-8), impact the immune response and pathogen clearance?
+6), impact the immune response and pathogen clearance?
 To isolate the effect of baseline propensities ($\vec{\ell}$) independently of signal feedback ($\vec{\psi}$), we first simulated “hard-wired” designs in which all signal sensitivities were set to zero ($\vec{\psi}_{i}=0$).
@@
 Interestingly, varying the baseline rate of differentiation $g_{0,N^{*}\to E}$ had a weaker impact on the overall effector profile and pathogen clearance (Fig.
-2B, C), suggesting relevance more downstream in the response program.
-Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated cytokine signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig.
+2B, C), suggesting its relevance may lie elsewhere in the response program.
+Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated (cytokine) signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig.
 S2).
@@
 Previous research has identified a trade-off between maximizing infection clearance and minimizing response toxicity (immunopathology) in the immune response [59, 28, 29].
-Consistently, by systematically sampling the space of designs (SI Appendix), we found this trade-off in the form of a Pareto front: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig.
+Consistently, by systematically sampling the space of designs (SI Appendix), we found the same trade-off: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig.
 3A).
@@
 This trade-off is stronger for infections with a higher intra-host reproductive number $R_{0}$ or lower immunogenicity $\kappa$.
-Interestingly, for scenarios with lower $R_{0}$, many of the hard-wired response programs without signal feedback ($\vec{\psi}_{i}=0$) lie along the Pareto front, and perform near-optimally.
-For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far inside the sub-optimal region.
+Interestingly, for scenarios with low $R_{0}$, many of the hard-wired response programs without signal feedback ($\vec{\psi}_{i}=0$; black dots) perform near-optimally, laying close to the Pareto front.
+Here, the Pareto fronts at the boundaries of these graphs indicate the set of strategies for which neither harm nor toxicity can be reduced without increasing the other, representing the best achievable balance between these competing objectives (see SI Appendix).
+For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far from the Pareto front and well inside the sub-optimal region.
 Proportionate control in these cases requires designs with appropriate sensitivities to antigen and cytokine/harm signals ($\vec{\psi}\neq 0$) to minimize infection harm while limiting immunopathology.
 Figure 4: Pareto-optimal designs balancing harm from acute infections and autoimmunity.
-(A) Bi-objective optimization to minimize the weighted sum of harms associated with T-cell regulatory designs $\vec{\theta}$ across two infection archetypes ($A,B$) yields a Pareto front (magenta curve); eq.
-10.
-Points on the curve are Pareto-optimal designs, parameterized by the weight $\lambda\in[0,1]$, quantifying the emphasis on archetype B; moving left to right corresponds to decreasing $\lambda$ (less emphasis on B).
-(B) The density plot shows harms incurred by $4.8\times 10^{6}$ uniformly sampled designs in their response to archetype I (non-replicating, self-like/ auto-immune inducing or allergenic antigens; $R_{0}=1$) versus archetype II (immunogenic infection; high $\kappa$), as defined in Fig.
+(A) Bi-objective optimization to minimize the weighted sum of harms associated with T-cell regulatory designs $\vec{\theta}$ across two pathogen archetypes ($A,B$) yields a Pareto front (magenta curve); eq.
+8.
+Points on the curve are Pareto-optimal designs, parameterized by the weight $\lambda\in[0,1]$, quantifying the emphasis on archetype A; moving left to right corresponds to decreasing $\lambda$ (less emphasis on archetype A).
+(B) The density plot shows harms (in units of the maximum number of susceptible cells $S_{\text{max}})$ incurred by $4.8\times 10^{6}$ uniformly sampled designs in their response to archetype I (immunogenic infection; high $\kappa$) versus archetype II (non-replicating, self-like/ auto-immune inducing or allergenic antigens; $R_{0}=1$), as defined in Fig.
 3B.
@@
 For each point $\lambda$ on the front, we identify the ensemble of designs closest to the Pareto front ($\Theta(\lambda)$), where a design belongs to the near-Pareto ensemble at $\lambda$ if its incurred harm against archetypes I and II simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of the respective archetypes.
-The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-I harm incur larger increases in archetype-II harm (see SI Appendix).
-The average design parameter values over the near-elbow ensemble are $\langle\vec{\theta}\rangle_{\Theta(\lambda^{*})}=(2.89,1.67,-0.67,-0.22,-1.94,0.94)$.
+The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-II harm incur larger increases in archetype-I harm (SI Appendix).
 (C) Decomposition of total harm along the Pareto front in (B) into infection-related harm and toxicity (eq.
-9).
+7).
 The black dashed line indicates the position of the elbow from (B).
-(D) Solid lines show parameter magnitudes for the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, as a function of harm incurred under archetype II along the Pareto curve (parameterized by $\lambda$).
-Shaded envelopes denote the standard deviation of the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front.
+(D) The magnitudes of the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, are shown as a function of harm incurred under archetype I along the Pareto curve (parameterized by $\lambda$).
+Solid and shaded envelopes show the mean and standard deviation of the design parameters, respectively, across the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front.
 (E) For each archetype in Fig.
-3B, total harm is shown for the archetype-specific minimizer $\vec{\theta}^{(\mathrm{min})}$ (circle) and for the average over near-elbow ensemble in (B).
-See SI Appendix, Figs.
-S6, S7 for Pareto fronts formed by other pairs of archetypes.
+3B, total harm is shown for the archetype-specific minimizer $\vec{\theta}^{(\mathrm{min})}$ (circle) and for the average over the near-elbow ensemble in (B).
+See SI Appendix, Fig.
+S6 for Pareto fronts formed by other pairs of archetypes.
 All simulation parameters other than the varied $K_{I}$, $b_{I}$, and $\vec{\theta}$ match Fig.
@@
 A design $\vec{\theta}$ belongs to this ensemble if its incurred total harm (eq.
-9) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection.
+7) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection; note that the maximum number of susceptible cells $S_{\text{max}}$ provides a convenient reference scale against which to quantify harm (cell death).
 Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, Fig.
 S3).
-Four coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig.
-3B): I—non-replicating, self-like/auto-immune inducing or allergenic antigens ($R_{0}=1$); II—highly immunogenic bacteria and viruses (high $\kappa$ or high $R_{0}$); III—intermediate replication and detectability, typifying mixed-strategy intracellular pathogens (mid-high $R_{0}$, mid-high $\kappa$); IV—difficult to detect threats such as growing tumors (low $\kappa$, broad range of $R_{0}$).
-The Pareto trade-off between infection harm and toxicity differs across these archetypes: designs optimized for archetypes I-III nearly eliminate infection damage with limited immunopathology, while designs in archetype IV show a strong trade-off (Fig.
+Three coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig.
+3B): I— highly immunogenic bacteria and viruses (high $\kappa$ or high $R_{0}$), associated with type-I immunity [60]; II— non-replicating, self-like/auto-immune inducing or allergenic antigens ($R_{0}=1$), in part associated with type-II immunity [60]; III— mid-range threats (low-medium $\kappa$ and $R_{0}$) including difficult to detect pathogens and slowly growing tumors.
+The balance between minimizing toxicity and clearing infection (i.e., reducing harm) differs across these archetypes, resulting in different design strategies: signal-sensitive designs optimized for archetypes I and III prioritize pathogen clearance at the cost of increased immunopathology, with the severity of the trade-off varying by scenario, whereas the largely pre-programmed designs optimized for archetype II suppress toxicity to avoid autoimmunity (Fig.
 3C).
@@
 3D).
-For archetype I—non-replicating antigens—the optimal design is largely pre-programmed to suppress the immune response: signal sensitivities $\vec{\psi}$ are close to 0, and the baseline rates are tuned to slow activation, proliferation, and differentiation, and enforce rapid effector deletion.
+For archetype II (non-replicating antigens) the optimal design is pre-programmed suppression of the immune response: signal sensitivities $\vec{\psi}$ are close to 0, and the baseline rates are tuned to slow activation, proliferation, and differentiation, and enforce rapid effector deletion.
 This damped response completely avoids toxicity when the threat does not grow so infection harm is determined by the initial exposure and cannot be decreased.
-Across all other archetypes a common pattern emerges—the response-derived cytokine signal $\sigma_{\text{res}}$ acts as an anti-inflammatory brake ($\psi_{\text{res}}<0$), with its magnitude tuned to context.
-Moreover, antigen sensitivity $\psi_{\text{Ag}}$ grows with immunogenicity: the effector program optimized for highly immunogenic pathogens in archetype II shows the largest $\psi_{\text{Ag}}$, whereas the programs for the poorly immunogenic archetype IV show diminished antigen sensitivity.
-Compared to archetype II, the baseline propensities $\ell$’s (or equivalently $g_{0}$’s) for archetype III accelerate effector contraction ($g_{0,E\to\emptyset}\,\uparrow$).
-For the weakly immunogenic archetype IV, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs.
+Across the other two archetypes a common pattern emerges: the response-derived cytokine signal $\sigma_{\text{res}}$ acts as an anti-inflammatory brake ($\psi_{\text{res}}<0$), with its magnitude tuned to context.
+Moreover, antigen sensitivity $\psi_{\rm Ag}$ grows with immunogenicity, with the designs optimized for archetype I (strong immunogenicity) showing the largest sensitivity to antigens, while those for archetype II (non-immunogenic) show no antigen sensitivity, and are largely hard-wired.
+Overall, across all archetypes, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs.
 S4, S5).
 Endogenous T-cell programs are shaped by minimizing harm from autoimmunity and acute infections
-Although the relative evolutionary pressures exerted by different threats are unknown, we can still pose a principled multi-objective problem for the effector program, i.e., find a design $\vec{\theta}$ that minimizes the weighted sum of harms across pathogen archetypes:
-$\vec{\theta}^{\prime}(\{\lambda_{i}\})=\underset{\vec{\theta}}{\arg\min}\sum_{i:\text{archetypes}}\lambda_{i}H^{(i)}(\vec{\theta}),$ (10)
+Although the relative evolutionary pressures exerted by different threats are unknown, we can still pose a principled multi-objective problem for the effector program, i.e., find a design $\vec{\theta}_{\text{min}}$ (eq.
+6) that minimizes the weighted sum of harms across a desired set of pathogen archetypes:
+$\vec{\theta}_{\text{min}}=\underset{\vec{\theta}}{\arg\min}\sum_{i:\text{archetypes}}\lambda_{i}H^{(i)}(\vec{\theta}),$ (8)
 where $\lambda_{i}\in[0,1]$ with the constraint $\sum_{i}\lambda_{i}=1$.
-The weights $\{\lambda_{i}\}$ determine the relative importance of each archetype in determining the optimal design $\vec{\theta}^{\prime}$.
-The solution to this family of problems forms a simplex over which $\lambda_{i}^{\prime}s$ vary (Fig.
+The weight $\lambda_{i}$ specifies how strongly we prioritize minimizing harm associated with the $i^{th}$ pathogen archetype (i.e., an objective), when identifying the optimal design $\vec{\theta}_{\text{min}}$.
+For $m$ archetypes, the solution to this family of problems maps an m1m-1-dimensional simplex of $\vec{\lambda}=(\lambda_{1},\dots,\lambda_{m})$ values to a set of designs that achieve the best possible trade-off between the different archetypes (Fig.
 4A; SI Appendix).
-This approach accommodates uncertainty about the empirical prevalence of each type of infection and the preferences/contingencies of evolution.
-Even though optimization can be done simultaneously over all four archetypes, we proceed by examining pairwise trade-offs that generate a one dimensional Pareto front, parameterized by $\lambda$ (Fig.
+When optimizing performance against one archetype necessarily worsens performance against another (i.e., a tradeoff), these optimal designs form a Pareto front: a set of solutions for which no objective can be improved without degrading at least one other (Fig.
 4A).
-A biologically salient case contrasts archetype I (non-replicating antigens) with archetype II (immunogenic pathogens).
-Non-replicating antigens plausibly arise from unprogrammed cell death events or external trauma inducing inflammation and may be implicated in autoimmune flares.
-This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Figs.
-S6, S7, for Pareto fronts formed by optimizing other pairs of clusters.
+In contrast, if there are no such conflicts among archetypes, the Pareto front collapses to a single design that is simultaneously optimal for all objectives.
+This framework naturally accommodates uncertainty in the prevalence of pathogenic challenges and in the effective “preferences” imposed by evolution or environment, in finding the optimal solution.
+Similar approaches have been successfully applied in other biological systems [61, 62].
+Even though optimization can be done simultaneously over all three archetypes, we proceed by examining pairwise trade-offs that generate a one dimensional Pareto front, parameterized by a single scalar $\lambda$ (Fig.
+4A).
+A biologically salient case contrasts archetype I (immunogenic pathogens) with archetype II (non-replicating antigens).
+Non-replicating antigens plausibly arise from un-programmed cell deaths or external trauma inducing inflammation and may be implicated in autoimmune flares.
+This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Fig.
+S6 for Pareto fronts formed by optimizing other pairs of clusters.
 Fig.
-4B shows harm from archetype I plotted against that of archetype II for a broad range of sampled designs and highlights the Pareto front (magenta).
+4B shows harm from archetype II plotted against that of archetype I for a broad range of sampled designs and highlights the Pareto front (magenta).
 Designs close to the Pareto front achieve near-best total protection for a given compromise weight $\lambda$; moving to the right along the Pareto front shifts the emphasis from clearance of immunogenic pathogens ($\lambda\uparrow$) to tolerance ($\lambda\downarrow$).
 Decomposing total harm into infection-related harm and toxicity (eq.
-9) shows that along this Pareto front it is relatively easy to reduce toxicity to non-replicating antigens without substantially worsening outcomes against immunogenic infections (Fig.
-4C).
+7) shows that along this Pareto front it is relatively easy to reduce toxicity to non-replicating antigens without substantially worsening outcomes against immunogenic infections; this is demonstrated in Fig.
+4C as substantial decreases in toxicity are achieved with only minor changes in infection harm.
 Figure 5: Elbow-optimal designs recapitulating macro-dynamics of T-cell expansion.
-(A,B) Fold expansion (i.e, peak clone size divided by the initial precursor number) of antigen-specific CD8+ T-cells in mice plotted against (A) the initial precursor number (data from [9]), and (B) the relative pMHC concentration required for half-maximal activation EC50 (data from [62]); in our model, EC50 is proxied by $K_{I}$.
-(C,D) Simulated mean fold expansion for designs belonging to the near-elbow ensemble from Fig.
-4B ($\Theta(\lambda^{*})$) when their sensitivity to the infection-harm signal ($\psi_{\text{inf}}$) is varied (colors), shown as a function of (C) the initial precursor T-cell number, and (D) the immunogenicity of the antigen $\kappa$.
+(A,B) In maroon (left vertical axis), fold expansion (i.e, peak clone size divided by the initial precursor number) of antigen-specific CD8+ T-cells in mice plotted against (A) the initial precursor number (data from [9]), and (B) the relative pMHC concentration required for half-maximal activation EC50 (data from [63]); in our model, EC50 is proxied by $K_{I}$.
+In blue (right vertical axis), simulated mean fold expansion for designs belonging to the near-elbow ensemble from Fig.
+4B ($\Theta(\lambda^{*})$) shown as a function of the initial precursor T-cell number (A), and the immunogenicity of the antigen $\kappa$ (B).
+Their sensitivity to the infection-harm signal ($\psi_{\text{inf}}$) is varied (shades of blue).
 For each design in the near-elbow ensemble identified in Fig.
-4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$ (colors), and simulate responses against the broad sample of infections (Fig.
+4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$, and simulate responses against the broad sample of infections (Fig.
 3B).
 Trend lines are obtained by averaging over infections and designs in the near-elbow ensemble (SI Appendix, Figs.
-S8, S9).
+S7, S8).
 Elbow designs with reduced sensitivity to infection-derived harm qualitatively match the experimental macro-dynamics of T-cell expansion.
 Unless otherwise indicated, simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}$.
-Table 1: Endogenous design of CD8+ T-cell response.
 Design Parameter Effect Source
@@
 6. response, $\psi_{\text{res}}$ suppress [29, 28]
-Next, we evaluated immune designs near the Pareto curve (Fig.
+Table 1: Endogenous design of CD8+ T-cell response.
+Next, we evaluated immune designs near the Pareto front (Fig.
 4D; SI Appendix).
 For each point $\lambda$, we identified an ensemble of designs near the Pareto front, $\Theta(\lambda)$.
-For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of those respective archetypes.
-The elbow of the Pareto curve represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype I will incur disproportionally large increases in harm from archetype II (see SI Appendix).
-The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$).
+For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the Pareto optimal harm of those respective archetypes at the same compromise value $\lambda$.
+The elbow of the Pareto front represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype II will incur disproportionally large increases in harm from archetype I (Fig.
+4B; SI Appendix).
+The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$); Fig.
+4D.
 Sensitivity to these signals diminishes as $\lambda$ decreases (i.e., as avoidance of autoimmunity is prioritized); see Fig.
@@
 4D).
-As $\lambda$ increases (leftward; clearance weighed more), all baseline rates shift toward promoting more effector cells.
-As $\lambda$ decreases below the elbow (rightward), the already slow differentiation baseline rate falls toward zero, while activation/proliferation and contraction are tuned to suppress effector production, reflecting designs that minimize collateral damage even at the expense of incomplete clearance.
-Notably, despite being optimized only for archetypes I and II, designs in the near-elbow ensemble $\Theta(\lambda^{*})$ perform competitively for other archetypes, incurring harm close to the archetype-specific optimal designs (Fig.
-4E; SI Appendix, Figs.
-S6, S7).
+As $\lambda$ increases (leftward; clearance weighes more), all baseline rates shift toward promoting more effector cells.
+As $\lambda$ decreases below the elbow (rightward; toxicity weighs more), the already slow differentiation baseline rate falls toward zero, while activation/proliferation and contraction are tuned to suppress effector production, reflecting designs that minimize collateral damage even at the expense of incomplete clearance.
+Notably, despite being optimized only for archetypes I and II, designs in the near-elbow ensemble $\Theta(\lambda^{*})$ perform competitively for archetype III, incurring harm close to the archetype-specific optimal designs (Fig.
+4E; SI Appendix, Fig.
+S6).
 From these analyses alone, it is not possible to ascertain the compromises encoded in the endogenous T-cell response program.
@@
 Figure 6: Engineering immune response for cancer immunotherapy.
-(A) The schematic shows the potential clearance of replicating cancer tumor cells by T-cells (e.g., infused during immunotherapy), and the potential collateral damage of therapy on susceptible tissue cells.
+(A) The schematic shows the potential clearance of replicating cancer tumor cells by T-cells (e.g., infused during immunotherapy), and the potential collateral damage of therapy to susceptible tissue cells.
 The accessible levers in an immunotherapy include the dosage and antigenicity of the infused cells.
@@
 Given a specific cognate T-cell, immunogenicity is the ratio of the antigenicity of a cancer cell to that of a healthy cell.
-(B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm as a function of cancer antigenicity and dosage of administered T-cells.
+(B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm (in units of the maximum number of susceptible cells $S_{\text{max}})$ as a function of cancer antigenicity and dosage of administered T-cells.
 Yellow points correspond to the optimal dosage of T-cells for a given immunogenicity/antigenicity.
-(C) Total harm, harm from cancer and toxicity as functions of cancer immunogenicity ($\kappa$) are shown at the optimal T-cell dosage (yellow circles in B).
+(C) Total harm, harm from cancer, immune toxicity, and the final tumor size following the immune response are shown as functions of cancer immunogenicity ($\kappa$) at the optimal T-cell dosage (yellow circles in B).
 (D) Clearance response—defined as the derivative of negative harm (i.e., protection) with respect to each design parameter $\theta_{i}$, $\partial H(\vec{\theta})/\partial{\theta_{i}}$ for $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$—averaged over the tuned near-elbow designs $\tilde{\Theta}(\lambda^{*})$, is shown for different levels of tumor immunogenicity $\kappa$; Left: $\vec{\psi}$ parameters; right: $\vec{g}_{0}$ parameters (colors).
@@
 This pattern was recapitulated by explicitly modeling competition for antigens in polyclonal T-cell responses [66].
-Near-elbow designs $\vec{\Theta}(\lambda^{*})$ (Fig.
-4) reproduce this inverse dependence on initial clone size but underestimate the steepness of the trend with respect to immunogenicity ($\kappa$) (Fig.
-5C, D).
-Reducing the sensitivity to infection harm $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamic range (Fig.
-5D) while having only a minor effect on total harm, and hence, our optimization objective (Fig.
-S9).
+Without any parameter fitting, near-elbow designs $\vec{\Theta}(\lambda^{*})$ from Fig.
+4 reproduce this inverse dependence on initial clone size but underestimate the steepness, and overestimate the magnitude of the trend with respect to immunogenicity ($\kappa$) (Fig.
+5A, B).
+Reducing the sensitivity to infection harm $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamics range (Fig.
+5B) while having only a minor effect on total harm, and hence, our optimization objective (Fig.
+S8).
 Interestingly, adjusting the other sensitivities ($\psi_{\text{Ag}},\psi_{\text{res}}$), or the baseline rates $g_{0}$’s does not reproduce the dynamic range observed in the experiments (Figs.
-S8, S9).
+S7, S8).
+Improved quantitative agreement may be achieved by adjusting a subset of the model’s non-design parameters (SI Appendix, Table S1), an analysis that would benefit from a larger dataset to enable robust fitting and validation; see SI Appendix and Figs.
+S10–S12 for the sensitivity of model performance to variation in non-design parameters and to simulation noise.
 To align responses with the observed macro-dynamic trends, subsequent analyses consider near-elbow designs $\vec{\theta}\in\Theta(\lambda^{*})$ with their infection-signal sensitivity reduced to one-tenth of their original value: for $\psi_{\text{inf}}\mapsto 0.1\,\psi_{\text{inf}}$.
 This ensemble of tuned designs is indicated by $\tilde{\Theta}(\lambda^{*})$.
-See SI Appendix, Figs.
-S11 and S12 for extensive model parameter sensitivity analyses with this ensemble.
 Designing an effector response to a growing tumor
@@
 6A; SI Appendix).
+We quantified harm as the total number of healthy cells lost, combining cell death caused directly by tumor expansion (replacing healthy cells) and collateral killing mediated by the immune response (Fig.
+6B,C).
 Similar to the case of infection control, a clearance-toxicity trade-off arises, but it is sharpened in cancer because many tumor antigens are only marginally distinct from self.
@@
 6B, C).
-As tumor immunogenicity increases, harm declines due to decreases in cross-reactive toxicity and increases in clearance (Fig.
+As tumor immunogenicity increases, harm decreases due to reduced cross-reactive toxicity and enhanced tumor clearance (Fig.
 6C).
@@
 TCR engineering raises an additional nuance: increasing TCR-tumor affinity often increases affinity to self [70], so that increasing tumor immunogenicity is difficult.
-To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged (SI Appendix, Fig.
-S10).
+To capture this, we evaluated treatment scenarios where the affinities/reactivities of tumor and self to the engineered TCR/CAR are proportional, leaving tumor immunogenicity unchanged ($\kappa=K_{I}^{-1}K_{S}=\text{constant}$; SI Appendix, Fig.
+S9).
 We quantify affinity or reactivity of a cell by the ratio of the maximum healthy cell counts $S_{\text{max}}$ to the number of target cells required for half-maximal T-cell response (EC50), i.e., $K_{I}^{-1}S_{\text{max}}$ for tumor cells, and $K_{S}^{-1}S_{\text{max}}$ for self targets.
-Provided that tumor immunogenicity ($\kappa=K_{I}^{-1}K_{S}$) is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig.
-S10A) mirrors its dependence on immunogenicity in Fig.
+Provided that tumor immunogenicity $\kappa$ is not too small and self-reactivity remains moderate, the dose-response as a function of tumor antigenicity (Fig.
+S9A) mirrors its dependence on immunogenicity in Fig.
 6B.
 Nevertheless, these cross-reactive regimes (Fig.
-S10B) yield higher total harm than scenarios where self-interaction fixed so that increasing tumor antigenicity increases immunogenicity (Fig.
+S9B) yield higher total harm than scenarios where self-interaction is fixed so that increasing tumor antigenicity increases immunogenicity (Fig.
 6C).
@@
 These include the extent to which affinity-driven TCR cross-reactivity can be mitigated (Fig.
-S10) and may require patient-specific measurements of T-cell interactions and signaling to calibrate designs for robust, durable therapy.
-Discussion
-In this work, we modeled the CD8+ T-cell response as a feedback-controlled program that integrates three cues—instantaneous antigen, cumulative infection-derived harm, and cumulative response-derived harm—to regulate activation, expansion, differentiation, and contraction of T-cells.
+S9) and may require patient-specific measurements of T-cell interactions and signaling to calibrate designs for robust, durable therapy.
+IV Discussion
+In this work, we modeled the CD8+ T-cell response as a feedback-controlled program that integrates three cues (i.e., instantaneous antigen, time-integrated infection-derived harm, and time-integrated response-derived harm) to minimize cumulative cell death during an immune challenge.
 Modulation of response proportional to the antigen signal and the time-integrated harm signals mirrors proportional-integral-derivative (PID) control of engineered systems, a strategy known for its stable and robust steering of dynamical processes toward desired outcomes [81].
 Exploring an ensemble of biologically motivated controllers (designs) and optimizing for pathogen clearance while minimizing immunopathology exposed a Pareto trade-off: a reduction in infection harm generally leads to an increased collateral damage to healthy tissues.
-This pattern persisted across diverse pathogenic challenges defined by their immunogenicity and within-host basic reproduction number.
-Three design principles emerged: (i) Pre-programmed (baseline-tuned) designs help restrain autoimmunity and slowly proliferating threats but are inefficient against fast-proliferating infections or metastatic tumors; effective control there requires signal integration and feedback.
+This pattern persisted across diverse immune challenges defined by their immunogenicity and within-host basic reproduction number.
+Three design principles emerged from our analyses: (i) Pre-programmed (baseline-tuned) designs help restrain autoimmunity and slowly proliferating threats but are inefficient against fast-proliferating infections or metastatic tumors; effective control there requires signal integration and feedback.
 (ii) Among signal sources, antigen and response-harm dominate; sensitivity to infection-harm contributes comparatively little once CD8+ effector T-cell expansion is underway.
-(iii) Feedback from response-harm acts as an anti-inflammatory brake, enabling strong antigen-driven responses while limiting collateral damage when the response overshoots—a characteristic consistent with endogenous effector programs.
-We posited that effector programs should attempt to solve a multi-objective problem to counter diverse pathogenic challenges.
+(iii) Feedback from response-harm acts as an anti-inflammatory brake, enabling strong antigen-driven responses while limiting collateral damage when the response overshoots—a characteristic consistent with endogenous effector programs [29, 28].
+Table 1 further summarizes the characteristics of our inferred optimal designs and relates them to supporting evidence from endogenous effector programs.
+We posited that effector programs should attempt to solve a multi-objective problem to minimize harm (cell death) inflicted by diverse pathogenic challenges.
 Interestingly, we found that a Pareto-optimal design that suppresses responses to frequent, self-like injuries while clearing recognizable, replicating (e.g., acute) infections mirrors the endogenous T-cell programs in vertebrates: strongly antigen-driven with an anti-inflammatory brake, and conservative cell-fate transition speeds that prevent runaway effector responses.
-We assumed that, although memory formation is essential, its immediate impact on primary effector magnitude is modest: $5-10\%$ of peak effector numbers [81, 16].
-Memory “quality” (longevity, proliferative capacity) depends on differentiation timing and the avidity of the selected clones [82, 15, 21, 83].
-To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall.
-An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the host’s evolutionary survival pressures that could have shaped the effector program.
+We note that, beyond minimizing harm due to cell death, additional factors may also shape the objectives optimized by effector T-cell programs, including the metabolic costs of effector activity and cytokine production, the cumulative burden of inflammation, and the long-term value of generating memory cells.
+Although the relative importance of these contributions remains unclear, they could be incorporated and analyzed within the same Pareto-optimization framework we introduced in this work.
+Memory formation is a key feature of CD8+ T-cell immunity, but here we focus on acute infection control and therefore do not model memory expansion or recall explicitly.
+In endogenous responses, immunological memory arises in part from the differentiation of a small fraction of effector cells ($\sim 5-10\%$ of the peak response [82, 16]) into effector-memory cells, effectively removing them from the active response.
+In our model, this loss of effector cells is captured by the contraction step, which can represent either effector apoptosis or differentiation into memory.
+By aggregating these processes into a single contraction term, we reduce degeneracy in the design parameters while retaining interpretability.
+Nonetheless, memory “quality” depends on differentiation timing and clone avidity [83, 15, 21, 84], which would be relevant for the strength and durability of protection upon secondary immune challenges.
+Extending our framework to an explicit allocation trade-off, where diverting cells to memory reduces primary effector output but seeds future protection, will be an important next step toward linking control designs to memory formation and the evolutionary pressures shaping effector programs.
 Using a minimal tumor model, we showed how the design objective informs rational immunotherapy protocols.
@@
 Second, treating the design parameters as engineering targets yields testable strategies to improve efficacy and safety, which is highly desirable [85, 86, 87].
-For moderately immunogenic tumors, promoting effector output by raising baseline rates of activation/proliferation or increasing antigen and infection sensitivities, while tempering the anti-inflammatory brake, can produce robust responses.
+For moderately immunogenic tumors, promoting effector output by raising baseline rates of activation/proliferation or increasing antigen and infection sensitivities, while tempering the anti-inflammatory brake, may produce robust responses.
 To operationalize these ideas, abstract signal sensitivities and baseline rates must be mapped to gene programs as targets of cell engineering (Table 1).
@@
 Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them?
+Would explicitly modeling interactions between CD8+ T-cells and other immune compartments, such as regulatory T-cells and innate immune system, enhance the robustness of cytotoxic responses across different pathogenic challenges?
 Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
-II Data and code availability
-Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune.
+V Materials and Methods
+Data and code for all the analyses can be found in the Github repository: https://github.com/StatPhysBio/designimmune. Previously published data from refs. [9, 63] were used for this work.
+Intra-host infection dynamics.
+We model a population of healthy, susceptible cells $S(t)$, initially consisting of $S_{\text{max}}$ identical cells.
+Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs.
+1A,B, S1).
+The dynamics of the susceptible and infected cells follow,
+$\begin{split}\dot{S}&=-b_{I}IS-d_{E}\frac{E}{K_{S}+I\frac{K_{S}+E}{K_{I}+E}+E+S}S\\ \dot{I}&=b_{I}IS-d_{I}I-d_{E}\frac{E}{K_{I}+S\frac{K_{I}+E}{K_{S}+E}+E+I}I,\end{split}$ (9)
+where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [92, 93, 94].
+The functional form of the last terms in these equations is motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix).
+Note that these equations describe dynamics during the immune challenge and exclude the slow birth-death processes that homeostatically maintain the susceptible cells in the absence of infection.
+Stochastic T-cell response model.
+We model the CD8+ T-cell response to an immune challenge as a time-inhomogeneous birth-death process coupled to deterministic dynamics of susceptible and infected cells.
+Beginning with an initial population of monoclonal naïve T-cells, we track individual T-cell lineages through recruitment of naïve cells to interact with antigen-presenting cells (APCs; $N\to\textbf{APC}\cdot N$), activation ($\textbf{APC}\cdot N\to N^{*}$), differentiation to effector ($N^{*}\to E$), expansion ($E\to 2E$), and contraction ($E\to\emptyset$).
+Naïve T-cells are recruited to APCs at a rate modulated by infection harm signals, and upon sustained engagement, become activated and undergo 2-3 burst-like divisions largely independent of continued antigenic input.
+During this early proliferative window, each activated cell commits to an effector or memory fate according to the first arrival of a time-inhomogeneous Poisson process with a signal-dependent rate; the number differentiating into effectors is drawn as a binomial random variable, with remaining cells adopting a memory phenotype.
+Effector cells subsequently divide and die at rates governed by time-varying antigen and harm signals.
+Proliferation is capped by a maximum division number [11, 7, 9, 10], which we set to 15 rounds, beyond which cells undergo terminal arrest (see SI Appendix for details).
+In our model, inter-lineage variability arises solely from stochastic differences in transition timing, capturing with prior observations [17, 18, 16, 19, 20, 15, 11, 21].
+To quantify this stochasticity and demonstrate robustness, results are averaged over five independent full-simulation replicates, with variability reported.
+The coefficient of variation averaged over immune challenges is small (C.V.
+= 0.05; Fig.
+S12), indicating that macroscopic quantities such as total harm and effector peak are robust to stochasticity in T-cell response dynamics.
+We also characterize sensitivity to simulation parameters (Figs.
+S10, S11), which vary across hosts and infection scenarios.
+Infection and response signals.
+T-cell activation relies on the integration of antigen and harm signals.
+Cell deaths can induce harm in the host and produce pro- or anti-inflammatory signals, depending on the underlying cause.
+We simplify this by distinguishing infection-induced and response-induced harms, which in a cumulative form, generate signals $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively.
+The instantaneous level of presented antigens, proxied by the infected-cell burden $I(t)$, defines a separate antigen signal $\sigma_{\text{Ag}}(t)$.
+Together, scaled by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
+$\begin{split}&\text{antigen signal:}\penalty 10000\ \penalty 10000\ \sigma_{\text{Ag}}(t)={K_{I}}^{-1}{I}(t)\\ &\text{harm signals:}\penalty 10000\ \penalty 10000\ \sigma_{\text{inf}/\text{res}}(t)=K_{H}^{-1}\int_{0}^{t}e^{-(t-s)d_{H}}\,h_{\text{inf}/\text{res}}(s)\,{\rm d}s.\end{split}$ (10)
+where $h_{\text{inf}}(s)$ is the number of cells killed by the pathogen (infection), and $h_{\text{res}}(s)$ denotes the number of cells killed by the effector T-cell response at time $s$ (eq.
+3).
+The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).
 Acknowledgement